### v1

# GoDe: Gaussians on Demand for Progressive Level of Detail and Scalable Compression

Abstract
3D Gaussian Splatting enhances real-time performance in novel view synthesis by representing scenes with mixtures of Gaussians and utilizing differentiable rasterization. However, it typically requires large storage capacity and high VRAM, demanding the design of effective pruning and compression techniques. Existing methods, while effective in some scenarios, struggle with scalability and fail to adapt models based on critical factors such as computing capabilities or bandwidth, requiring to re-train the model under different configurations. In this work, we propose a novel, model-agnostic technique that organizes Gaussians into several hierarchical layers, enabling progressive Level of Detail (LoD) strategy. This method, combined with recent approach of compression of 3DGS, allows a single model to instantly scale across several compression ratios, with minimal to none impact to quality compared to a single non-scalable model and without requiring re-training. We validate our approach on typical datasets and benchmarks, showcasing low distortion and substantial gains in terms of scalability and adaptability. The source code, provided in the supplementary material, will be open-sourced upon acceptance of the paper.
1 Introduction
Neural Radiance Fields (NeRFs) [25] have consistently boosted novel view synthesis, achieving state-of-the-art quality in scene reconstruction. However, its limitations lie in slow training and inference times. Over the last few years, different methods [13, 33, 5, 27, 12, 16] have been proposed to enhance speed, but none have successfully achieved real-time performance for complex, high-resolution scenes. 3D Gaussian Splatting (3DGS) [17] represents a major advancement by modeling scenes as mixtures of Gaussians, effectively replacing the computationally intensive tasks of ray casting, ray sampling, and neural network queries. By leveraging differentiable rasterization and spherical harmonics for handling view-dependent effects, 3DGS achieves real-time performance even in complex real-world scenarios and at high-definition resolution.
Figure 1: Our method constructs a progressive hierarchy of Gaussians, enabling progressive LoD. Combined with compression techniques, this allows the model to operate across several (eight in the figure) compression rates, ranging from extremely high to low. The model can dynamically adjust the rate based on criteria such as a target minimum FPS.
3DGS offers competitive training times and reconstruction quality that can rival or even surpass that of NeRF-based models. Despite these advancements, some challenges still persist. Specifically, 3DGS models often require large storage and VRAM resources, as scene representations may involve millions of Gaussians. This may hinder real-time rendering on consumer-level devices and make on-demand streaming of 3DGS model quite prohibitive. This has sparked interest in pruning and compression techniques to reduce VRAM consumption during runtime and minimize storage requirements, with high rendering quality. While current pruning and compression methods [2, 1, 29, 14, 37, 11] have shown effectiveness in improving performance and reducing VRAM usage, they lack scalability: they require further optimization and personalization for each specific compression level, resulting in multiple models, one for each compression level.
To address this scalability issue, we propose GoDe (Gaussians on Demand), a novel approach that enables scalable compression. As shown in Fig. 1, GoDe enables a 3DGS model to scale across different compression rates, without the need for retraining specialized models. Specifically, GoDe employs a progressive Level of Detail (LoD) approach: Gaussians are organized into discrete and progressive levels, with each level adding more detail (enhancement) to the previous one. This progressive structure allows 3DGS models to be highly scalable across different scenarios. In fact, by adding or removing enhancement layers, we can easily adjust the trade-off between quality and performance. When combined with compression technique schemes, this approach enables scalable compression. Each level of detail can be seen as a distinct compression level, where higher LoDs corresponds to lower compression rate and lower LoDs corresponds to a higher compression rate. This allows for navigating the rate/distortion curve by simply adding or removing enhancement layers. To build the hierarchy, inspired by iterative gradient-informed pruning approaches, we developed an iterative gradient-informed masking methods that progressively constructs the hierarchy in a top-down manner (from high to low levels). This methods is general and compatible with popular 3DGS baselines. A fine-tuning phase ensures coherence and smoothness across the levels, while a final compression stage further reduce model size.
In summary, our main contributions are:
  • A general method for progressive LoD (Sec. 3.1): Leveraging a novel gradient-based iterative masking technique, we show the applicability of this method to a variety of popular 3DGS model (Sec. 4.2) and how this approach effectively organizes primitives into a hierarchy of progressive levels with respect to other approaches (Sec. 4.4.2).
  • Smooth transition between levels (Sec. 3.2): Our fine-tuning procedure offers a simple and effective way to reduce artifacts between levels, ensuring smooth transitions in visual fidelity across multiple LoDs (Sec. 4.3) and boosting LoD quality (Sect. 4.4.3).
  • Scalable Compression (Sec. 3.3): Our method integrates and extends state-of-the-art compression techniques, enabling scalable compression and adaptive rendering, with comparable and sometimes even superior performance of non-scalable methods (Sec. 4.4.4).
We extensively evaluate our approach on standard benchmark and compared it against the most recent state-of-the-art methods, showing comparable and even superior performances than non-scalable method. Our method allows models to scale across a wide range of compression levels, from low to very high, achieving up to 99.76% reduction in model size and 98.36% in the number of primitives in the highest compression rate with respect to non-compressed models, while preserving comparable full-size model quality at the lowest compression rate.
2 Background and Related Work
2.1 3DGS Preliminaries
3DGS [17] has emerged as a cutting-edge technique for real-time radiance field rendering, delivering top-tier performance in both quality and speed. It represents a scene using volumetric Gaussian primitives, each defined by a set of parameters: a position $\mathbf{x}$, a covariance matrix $\mathbf{\Sigma}$, which is decomposed into scaling and rotation components, an opacity $o$, and Spherical Harmonics (SH) coefficients to capture view-dependent color or appearance. In the rendering stage, the 3D primitives are projected into 2D screen space and rasterized with alpha blending $\alpha=oG$ [39] with the projected Gaussian’s contribution to position $\mathbf{x}$ given by $G(\mathbf{x})=\exp{\left(-\frac{1}{2}\left(\mathbf{x}^{T}\right)\Sigma^{-1}\left(\mathbf{x}^{T}\right)\right)}$. The combined effect of converting SHs to per-view color values and applying alpha blending reconstructs the appearance of the captured scene. These primitives originate from a sparse Structure-from-Motion (SfM) point cloud and are optimized through a gradient-descent process, combined with adaptive refinements. This ensures the optimization of both pixel-wise similarity and structural similarity.
2.2 3DGS Level of Detail
LoD is a well-established technique in computer graphics, used to optimize performance and improve visual quality. It comprises three main components: constructing a scene representation at various levels of detail, implementing a policy for selecting the appropriate level based on the view, and enabling seamless transitions between levels to avoid artifacts or noticeable stuttering. Recent works have integrated LoD techniques with 3DGS, mainly to reduce aliasing artifacts [38] and improve performance, especially for large-scale datasets [23]. These methods typically involve learning multiple levels of detail within the scene representation, followed by selective rendering according to heuristics like pixel coverage (i.e., the number of pixels occupied by a Gaussian projection on the image plane) and distance. Other work propose a hierarchical representation, often structured as a tree, as proposed by [24, 18]. This method dynamically implements LoD by selectively rendering Gaussians up to a specific level within the structure based on certain heuristics. In these approaches, LoD is typically defined at render time (dynamic LoD), which complicates its integration into compression contexts like ours, where a predefined, discrete set of progressive enhancement layers is essential. Additionally, these dynamic LoD methods often introduce significant overhead due to the need for tree traversal algorithms to construct levels of detail, as opposed to the simpler and faster operation of layer concatenation, which is more suitable for adaptive rendering in bandwidth-limited scenarios.
2.3 3DGS Compression
Since the introduction of hybrid radiance fields models that incorporate auxiliary data structures to accelerate training and inference times, such as voxel grids [33, 5, 13, 28], research has focused on compression [8] and complexity reduction [9]. For instance, [8] demonstrates how an iterative-magnitude pruning approach is highly effective in reducing the number of parameters even in the context of radiance fields. These techniques have recently been adapted in the field of 3DGS. In fact, despite its advantages, 3DGS is limited by high storage and memory demands, reducing its application, for example, on low-end hardware. Current compression methods aim at achieving the highest quality at minimal storage size. Typically, compact representations involve post-hoc Gaussian pruning followed by compression stages, including quantization and entropy coding [1, 10, 30, 36]. Some methods operate directly during training, aiming to limit the number of Gaussians via learned masks and optimized encoding schemes [14, 6, 22, 31, 34]. Other studies enhance the representation by employing neural networks for on-the-fly Gaussian generation [24], tri-planes [35], or by using auxiliary structures such as hash grids to store spherical harmonics [6, 21], further reducing storage demands. In particular, Scaffold-GS [24] reduces storage by up to 5 times by replacing Gaussians with anchor points. These anchor points, which are significantly fewer in number, are associated with a set of features, offsets, and scalings. A set of shallow neural networks process each anchor point and generates the parameters of view-adaptive Gaussians. However, it is important to note that none of these methods are scalable in practice. While different compression rates can be achieved, they require retraining the model from scratch or performing additional fine-tuning steps for each compression level. This limitation makes it difficult to adapt these model across a broad range of context, such as low bandwidth or varying hardware capabilities. In contrast, our method incorporates a scalable compression scheme: with a single training we can instantly obtain different compression levels, effectively navigating the rate-distortion curve. Furthermore, our method matches or even surpass dedicated, non-scalable compression techniques.
3 Gaussians on demand
Figure 2: Graphical overview of our method GoDe. 1) Starting from a pre-trained model, we build a progressive hierarchy of Gaussians through iterative gradient-informed masking. 2) A fine-tuning stage further enhances the model’s quality and reduces its memory footprint by culling unnecessary SH coefficients. 3) The model is further compressed through codebook quantization.
This section presents our proposed efficient and adaptive 3D rendering method using hierarchical Gaussian representations. The proposed approach extends the standard 3D Gaussian Splatting (3DGS) framework by introducing a hierarchical representation of the scene at multiple levels of detail. The objective is to develop a single, scalable model that can dynamically adapt to various criteria, such as computational resources or available bandwidth. We present a novel algorithm that implicitly constructs this hierarchy, enabling a single 3DGS model to learn multiple levels of detail. We first describe how to build a progressive hierarchy of Gaussians from a pre-trained 3DGS model (Sec. 3.1) . Then, we show how to train and fine-tune Gaussians in different levels of detail (Sec. 3.2). Finally, we describe our scalable compression and adaptive rendering steps (Sec. 3.3).
3.1 Hierarchy of Gaussians
Let $L$ denote the total number of levels of detail in the hierarchy. To implement progressive LoD, our objective is to organize the Gaussians of model $\mathcal{G}$ as a sequence of progressive levels $l\in\{0,\dots,L-1\}$ defined as
$\mathcal{G}=(\mathcal{G}_{0},~{}\mathcal{E}_{1},\dots,\mathcal{E}_{l},\dots,\mathcal{E}_{L-1}),$ (1)
where $\mathcal{G}_{0}$ denotes a set of starting Gaussians, and $\mathcal{E}_{l}$ is the set of enhancement Gaussians at level $l$. We also refer to $\mathcal{G}$ as $\mathcal{G}_{L-1}$. The progressive LoD $\mathcal{G}_{l}$ at a given level $l$ is defined as:
$\mathcal{G}_{l}=\mathcal{G}_{0}\cup\bigcup_{i=1}^{l}\mathcal{E}_{i}.$ (2)
The sets are organized progressively, ensuring that
$\mathcal{G}_{0}\cap\mathcal{E}_{1},\dots,\cap~{}\mathcal{E}_{l},\dots,\cap~{}\mathcal{E}_{L-1}=\emptyset.$ (3)
This guarantees that each level of detail progressively introduces new Gaussians, with no overlap among the layers. In order to define how many Gaussians are associated with each enhancement set, we define the number of Gaussians $|\mathcal{G}_{l}|$ associated with level of detail $l$ as:
$|\mathcal{G}_{l}|=c_{\text{min}}\cdot b^{l},$ (4)
where $b$ is computed as
$b=\exp\left(\frac{\log(|\mathcal{G}_{L-1}|)-\log(c_{\text{min}})}{L-1}\right)$ (5)
and $c_{\text{min}}=|\mathcal{G}_{0}|$ is the number of Gaussians at the first LoD level (chosen as desired). We choose this progression since the rate-distortion functions generally follow a logarithmic law (see the theoretical rate-distortion curve for a Gaussian source [7]).
To organize the Gaussians of the model into a hierarchical structure, we employ an iterative gradient-informed masking process. This approach follows a top-down strategy, starting with the construction of the higher enhancement levels first, namely, progressively building the model’s hierarchy from layer $l=L-1$ to layer $l=1$. The gradient of each parameter in the Gaussians serves as a critical and general indicator of its relevance to the overall model. A Gaussian with a low gradient indicates minimal contribution to the loss function. Conversely, a high gradient suggests that the Gaussian is essential and has a significant impact on the loss. We can then leverage gradient information to build a hierarchy. Specifically, to create the enhancement set $\mathcal{E}_{l}$, we accumulate gradients on $\mathcal{G}_{l}$ using the training set in order to select the $k_{l}$ Gaussians with the lowest importance score defined as the sum of the norm of the gradient of all parameters of each Gaussian:
$\mathcal{E}_{l}=\texttt{lowest}_{k_{l}}\left\{\left\|\frac{\partial\mathcal{L}}{\partial\mathbf{\theta}_{i}}\right\|_{2}\right\}_{i\in\mathcal{G}_{l}},$ (6)
where, with some abuse of notation, $\texttt{lowest}_{k_{l}}$ denotes the set of the $k_{l}$ Gaussians with the lowest importance. Note that the difference $|\mathcal{G}_{l}|-|\mathcal{G}_{l-1}|$ is our $k_{l}$. The selected Gaussians will be associated with $\mathcal{E}_{l}$ and excluded (masked) from the next iterations. We reset the gradients and repeat the process. At the end of this iterative process, important Gaussians (with high gradient values) are placed at lower layers, while less important Gaussians are assigned to higher levels, where they model finer details and geometries. Although this phase requires $L-1$ steps, it tends to be very fast (1-2 minutes in total to build the hierarchy) as it only requires a single pass through the training set per step.
3.2 Fine-Tuning and SHs culling
This phase involves $F$ fine-tuning iterations designed to improve reconstruction quality and enforce a smooth transition between levels of detail. A random level $l$ is uniformly sampled at each iteration and a fine-tuning iteration is performed only on the $\mathcal{G}_{l}$ Gaussians associated with the selected level $l$. This optimization further improves reconstruction quality, especially at the lower levels. During fine-tuning, Gaussian parameters are reduced by selectively culling the spherical harmonics. Following the approach proposed by [31], SH coefficients are analyzed halfway through the fine-tuning phase to identify which primitives can be represented with 0, 1, 2, or 3 bands. The SH function is evaluated across all input views to determine whether a primitive requires view-dependent effects, based on the extent of color variation between views. If a primitive maintains a consistent color across all views, view-dependent effects are considered unnecessary. In addition, higher-order bands can be discarded if the color remains stable with lower-order SH coefficients. This phase significantly reduces the model’s memory cost (with negligible quality loss) since each Gaussian is represented by 59 32-bit floating point numbers, of which 45 are dedicated to the SH coefficients.
3.3 Scalable Compression
After fine-tuning, the compression step is performed. Compression is applied independently to the first LoD $\mathcal{G}_{0}$ and the following enhancements $\mathcal{E}_{l}$. Initially, a codebook is generated for each parameter (excluding the position of the primitives) using $K$-means clustering. In particular, following [31], codebooks of 256 entries are created for the opacity, the three scaling components, the real and imaginary parts of the rotation quaternion, the base color coefficients, and one for the 3-channel color component for each of the 15 SH coefficient groups. Codebook entries and positions are further quantized to 16-bit half-precision floats: in this way, we have achieved scalability since every LoD can be progressively decoded starting from the base one. It is worth pointing out that we move along the rate/distortion curve by just concatenating further enhancements. As shown in Sec. 4, our results are very competitive with the recent state-of-the-art, with the added value of a single scalable model obtained with a single training process. In conclusion, GoDe achieves both scalable compression and progressive LoD. Scalable compression can be exploited in many application scenarios beyond the simple memory-saving result, e.g. adaptive HTTP streaming [32]. The hierarchical organization of GoDe enhancements can be exploited for adaptive rendering at different LoD: specifically, the enhancements can be progressively splatted until a given target criterion, e.g. minimum frame rendering rate, is met. If rendering speed is the only goal, the GoDe compression phase can be skipped.
4 Experiments
4.1 Setup
The code was developed in Python 3.9 using PyTorch 2.1 and executed on a single NVIDIA A40 GPU. We set $L=8$ for all the configurations, $c_{\text{min}}=100,000$. The fine-tuning phase comprises $F=30,000$ iterations. All the other parameters remain unchanged from the original implementation. Our method was tested with three popular 3D Gaussian Splatting models: 3DGS [17], MCMC-3DGS [19], a variant of 3DGS that treats 3D Gaussians as samples from a Markov Chain Monte Carlo process, generally yielding slightly better results than 3DGS, and the neural-based Scaffold-GS [24]. For Scaffold-GS, due to its unique architecture, we adopted a slightly different compression scheme. This approach is similar to our proposed GoDe method but excludes SHs culling since SHs are not present in the model. Instead, it incorporates iterative masking with quantization-aware fine-tuning and entropy coding. The iterative masking process follows a similar approach; however, rather than masking Gaussians, we mask anchor points directly, evaluating the gradients of their respective features. Quantization is performed during the fine-tuning phase: features are quantized to 8-bit unsigned integers, while the scaling, offset, and positional attributes are quantized to half-precision floating points similarly to [30] After fine-tuning, the model parameters are further compressed using ZIP, which combines LZ77 compression with Huffman entropy coding to achieve efficient and lossless storage reduction.
Table 1: Comparison of our methods against recent methods. All methods have been retrained and tested on the same hardware. For clarity, we present four compression levels out of 8 for our configurations, selected to match the compression rate of the other methods. These levels, labeled as Ours, are derived from the same model without the need for retraining (in contrast to the other approaches). Red, orange and yellow respectively indicates the first, second and third best result.
Method Mip-NeRF360 Tanks&Temples Deep Blending
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size[MB] $\downarrow$ W $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size[MB] $\downarrow$ W $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size[MB] $\downarrow$ W $\downarrow$ FPS $\uparrow$
HAC (low)  [6] 27.78 0.812 0.277 22.0 502 68 24.37 0.853 0.215 11.1 307 88 30.34 0.909 0.329 6.3 190 166
HAC (high)  [6] 27.53 0.807 0.290 15.2 426 72 24.05 0.846 0.222 7.9 300 85 29.98 0.904 0.340 4.1 188 173
RDO  [34] 27.05 0.801 0.288 23.3 1860 178 23.32 0.839 0.232 11.9 912 257 29.72 0.906 0.318 18.0 1478 200
LightGaussian  [10] 27.24 0.810 0.273 51.0 2197 161 23.55 0.839 0.235 28.5 1210 225 29.41 0.904 0.329 43.2 956 348
SOG  [26] 27.08 0.799 0.277 40.3 2176 132 23.56 0.837 0.221 22.8 1242 227 29.26 0.894 0.336 17.7 890 236
Compact-3DGS   [21] 27.02 0.800 0.287 29.1 1429 108 23.41 0.836 0.238 20.9 842 147 29.76 0.905 0.324 23.8 1053 144
Reduced-3DGS [31] 27.19 0.810 0.267 29.3 1436 234 23.55 0.843 0.223 14.0 656 358 29.70 0.907 0.315 18.3 990 304
EAGLES [14] 27.13 0.809 0.278 57.1 1290 156 23.20 0.837 0.241 28.9 651 227 29.75 0.910 0.318 52.4 1192 144
Comp-GS [22] 27.28 0.802 0.283 16.4 485 165 23.62 0.837 0.248 9.8 241 286 29.89 0.904 0.336 10.0 246 253
27.75 0.810 0.284 19.7 453 98 24.01 0.846 0.226 11.0 266 142 30.34 0.909 0.331 8.4 63 212
27.59 0.803 0.297 14.1 369 107 23.95 0.843 0.234 8.9 230 156 30.29 0.908 0.334 7.2 56 208
27.34 0.793 0.316 10.3 288 127 23.85 0.837 0.244 7.3 196 176 30.14 0.905 0.343 5.3 44 225
Ours w/ Scaffold-GS 27.07 0.780 0.336 7.5 221 146 23.72 0.830 0.258 5.9 165 189 29.76 0.897 0.358 3.9 33 236
27.42 0.815 0.263 43.9 1546 146 23.97 0.842 0.220 22.2 938 216 29.71 0.901 0.323 19.5 931 263
27.23 0.804 0.289 20.0 603 222 23.76 0.831 0.241 11.8 438 311 29.70 0.901 0.326 11.0 491 345
26.99 0.790 0.312 13.4 381 268 23.49 0.821 0.259 8.6 301 366 29.66 0.901 0.331 8.3 357 391
Ours w/ 3DGS-MCMC 26.59 0.772 0.339 9.0 242 317 23.06 0.806 0.281 6.3 208 426 29.56 0.897 0.348 4.8 189 488
27.27 0.807 0.273 40.5 1546 189.1 23.76 0.839 0.231 21.1 938 269 29.73 0.904 0.327 19.8 931 302
27.16 0.801 0.295 18.3 603 270 23.66 0.832 0.245 11.6 438 383 29.73 0.903 0.334 10.1 491 425
26.93 0.791 0.315 12.0 381 317 23.48 0.824 0.259 8.6 301 439 29.74 0.902 0.340 7.2 357 492
Ours w/ 3DGS 26.55 0.773 0.341 7.9 242 371 23.13 0.812 0.279 6.4 208 509 29.67 0.899 0.348 5.1 189 564
4.2 Quantitative Results
Figure 3: Rate-distortion curves (left), PSNR/W (center, where W is the number of primitives) and PSNR/FPS (right) curve on Mip-NeRF360. Our approach achieves competitive results with both neural-based Gaussian models and explicit ones while maintaining significantly higher speed at comparable quality. It is also the only method that achieves scalability without requiring ad-hoc training.
We evaluated our method on the same datasets used by 3DGS, namely: Mip-NeRF360 [3], TanksAndTemples [20], and DeepBlending [15]. For each test, we report image quality metrics (PSNR, SSIM, LPIPS, computed on VGG), the model size in megabytes (SIZE), the number of primitives $|\mathcal{G}_{L-1}|$ (here $W$ for brevity), and the average frame rate on the test set (FPS). We present the results for each architecture, comparing them with various state-of-the-art proposals. To ensure a fair comparison, all the methods were re-trained and evaluated on our machine, for known discrepancies in reported results as noted by [4]. Additionally, some methods were re-trained using updated hyperparameters to generate rate-distortion curves in alignment with the parameters specified in their respective papers. The main results are summarized in Tab. 1. In Fig. 3, we plot rate-distortion and PSNR/FPS curves. These findings show that our method is generally competitive with current state-of-the-art approaches while benefiting from being a scalable model, as our rate/distortion curve is generated instantly. We divide our analysis of the results between neural-based 3DGS models and explicit ones, as the former generally provides superior rate-distortion performance.
Neural-based 3DGS
To the best of our knowledge, HAC [6] represents the current state-of-the-art in terms of rate distortion. HAC uses the Scaffold-GS [24] codebase. Scaffold-GS provides slightly higher reconstruction quality than 3DGS (approximately +0.3 dB) while reducing storage by about 5$\times$. By applying our method to Scaffold-GS we observe that it delivers comparable results, slightly outperforming HAC on the MipNeRF-360 dataset, while HAC remains superior on the Tanks&Temples dataset. We consider our results competitive, as our scalable compression method does not require re-training, unlike other methods that need to train a separate model with different hyperparameters. Our method offers another advantage beyond scalability: by masking a higher number of anchor points on average, our model is faster, achieving a frame rate increase ranging from +33% to +62% while maintaining similar reconstruction quality.
Explicit 3DGS
Regarding the comparison with explicit Gaussian splatting models, our results are state-of-the-art on nearly every dataset in terms of rate/distortion and performance at equivalent or even superior quality, achieving up to a 3$\times$ increase in FPS. There is one point where Comp-GS [22] slightly outperforms our method, but this approach scales less effectively at higher compression rates. In this case, our approach also offers the advantage of utilizing a single scalable model, which eliminates the need for re-training when adapting to different compression rates. Notably, the models obtained with our method are highly scalable, ranging from an extremely high compression rate of a few megabytes to an extremely low one.
4.3 Visualizing LoDs
In Fig. 4 we visualize some LoD levels for various scenes. For each LoD level, we report the average frame rate computed for rendering the image of the test set, the number of Gaussians, the average PSNR, and the model size. The transition between LoDs is smooth thanks to the fine-tuning process with random LoD sampling. Moreover, we observe that lower LODs are extremely fast and small, with rendering speeds exceeding 800 FPS and model sizes as low as 2.4MB for the drjohnson scene from the DeepBlending dataset (in comparison with the vanilla-3DGS which renders the same scene at 146FPS and requires 780MB of storage). Similar performances are achieved across other datasets. It is important to note that lower LoDs tend to lose fine details, particularly in high-frequency areas such as reflections, transparency handling, and background elements. This is particulary noticeable in the kitchen scene, where at lower levels of detail the glass bottle in the background is either not rendered at all or only partially and becomes progressively more defined at higher levels. Another example can be seen in the counter scene, with the reflection on the dishes. Other examples are the paining in the drjohnson scene and the trees in the background of the $truck$ scene. This is a significant result, demonstrating the effectiveness of the method, which automatically learns a hierarchy and preserves the most important details at each level, minimizing the loss of detail to secondary aspects. This is further illustrated in Fig. 5, where we render only the Gaussian associated with an enhancement level. As the level increases, we observe that the Gaussians focus on more complex geometry and finer details.
Figure 4: Visualization of LoDs. For visual purposes, we show the first three LoD levels and the last one. For each LoD we provide the average FPS, PSNR, primitives count W (in thousands), and size. Starting from top to bottom we have: kitchen, counter, drjohnson and truck scenes.
Figure 5: Visualization of enhancement levels. We render only the Gaussians belonging to the specified enhancement level.
4.4 Ablations
Here, we investigate how different choices of parameters and algorithms affect the render quality. These choices include the number of levels $L$, the ranking methods used to build the masks, the impact of compression, and the hierarchy on quality.
4.4.1 Number of levels
In our configuration, we opted for $L=8$ levels. However, we tested our method with a variable number of levels and observed minimal differences in performance. Fig. 8 shows that, with increasing number of levels, there is a slight decrease in the quality of lower LoDs, while the quality in higher levels remains similar. We chose 8 levels as a good balance, but our method is easily extendable to a higher number of levels without significant loss in quality. This highlights the scalability of our approach.
4.4.2 Pruning methods
In Fig. 8, we investigate multiple strategies for ranking Gaussians, which are commonly used to prune and compact the model. We found that our method provides better reconstruction quality with the same number of Gaussians compared to several proposed strategies, such as those based on opacity [17] or positional gradient [1]. Additionally, we compared our technique with the one proposed by LightGaussian [10] and still achieved superior results. All these methods have been tested using the same iterative process. We also evaluated the necessity of iterative masking versus the same ranking method but applied in a one-shot manner labeled as Ours (One Shot), where the gradient is accumulated only once to create the levels. Our experiments demonstrate the importance of the iterative approach, as the one-shot method leads to a loss in quality, particularly at lower LoDs. Since the iterative masking process is very fast, we opted to use the sum of the gradient norms of the parameters as a score and perform the masking iteratively.
Figure 6: Ablation on the number of levels on Mip-NeRF360.
Figure 7: Comparison among pruning methods on Mip-NeRF 360.
Figure 8: Comparison among our method with Scaffold-GS, a non-scalable variant of the same method and HAC. This constitutes an upper bound to our scalable method.
Table 2: Effects of fine-tuning and compression on Mip-NeRF360.
LoD W Mask Only Fine Tuning Fine-Tuning + Compression
[…]
0 100 18.52 0.553 24.1 25.20 0.719 24.1 25.24 0.717 3.5
1 155 19.90 0.606 36.7 26.06 0.752 36.7 26.03 0.749 5.3
2 242 21.52 0.662 56.4 26.65 0.778 56.4 26.55 0.773 7.9
3 381 23.28 0.715 89.1 26.84 0.787 89.1 26.93 0.791 12.0
4 603 25.00 0.760 142.8 27.08 0.796 142.8 27.16 0.801 18.3
5 962 26.45 0.793 216.6 27.33 0.807 69.8 27.25 0.806 27.5
6 1546 27.23 0.810 365.9 27.45 0.813 227.2 27.27 0.807 40.5
7 2497 27.49 0.815 562.2 27.47 0.814 365.6 27.29 0.808 57.9
AVG 811 23.67 0.714 186.7 26.76 0.783 126.5 26.72 0.781 21.6
Figure 9: Effects of fine-tuning on the train scene of Tanks&Temples dataset. In the first row, LoDs without fine-tuning; on the second, LoDs after fine-tuning.
4.4.3 Effects of Fine-Tuning and Compression
We analyze the effects of fine-tuning and compression from both a quantitative and qualitative perspective. As shown in Tab. 2, the fine-tuning phase significantly improves reconstruction quality, particularly for lower LoDs, while slightly reducing quality in the two highest LoDs, of about 0.02 dB. Regarding compression, we observe a reduction in storage ranging from $99.37\%$ to $89.69\%$, depending on the level considered. The impact on reconstruction quality is minimal, with a loss of 0.03 dB to 0.18 dB across different levels. In Fig. 9, we also visualize the visual impact of fine-tuning. Non-fine-tuned models exhibit holes and generally lower quality, while the LoDs after fine-tuning show improved quality with way fewer artifacts.
4.4.4 Impact of hierarchical representation
In this ablation study, we evaluate our hierarchical compression method against a non-hierarchical variant to assess how much our hierarchical approach diverges from the best achievable performance. Specifically, we re-run our method but omit any hierarchy constraints. The process is as follows: we accumulate gradients for one epoch, mask the $k$ Gaussians with the lowest gradient values, perform a fine-tuning cycle of $7,000$ iterations, and evaluate the performance (each evaluation is a point in Fig. 8). This procedure is repeated across 8 levels, down to a minimum threshold of $100,000$ parameters, with each level acting as an independent model. Interestingly, our method sets a new performance benchmark, outperforming HAC and underline the effectiveness of iterative, gradient-based pruning for this task. Our hierarchical version, though also superior to HAC, performs less well, particularly at high compression levels. This is due to the imposed hierarchy, which requires optimizing each level in relation to the others, thus introducing a trade-off between higher and lower levels.
5 Limitation and Future Work
Our approach demonstrates effectiveness in scalable compression tasks, matching or even surpassing the performance of non-scalable compression methods. However, it also inherits certain limitations typical of top-down pruning and compression approaches, such as the requirement for a pre-trained model with a well-established geometric prior. In future work, we intend to explore the potential for a bottom-up approach, integrating the hierarchical structure during the training of the Gaussian model.
6 Conclusion
This paper presented GoDe, a framework for a progressive level of detail and scalable compression. Our extensive experiments show that our method is highly scalable and competitive, often surpassing non-scalable state-of-the-art techniques in both quality and performance. In fact, unlike previously proposed methods, our approach allows for model compression at multiple rates without the need for retraining with different hyperparameters. This framework also enables adaptive rendering: given the hierarchical-layered structure of the model, it is possible to dynamically select Gaussians up to a certain level to meet specific criteria, such as a target frame rate. Additionally, a notable advantage of our approach is its generality: relying on widely applicable techniques, such as gradient-informed pruning, our method can be integrated into popular 3DGS models.

### v2

@@
 3D Gaussian Splatting enhances real-time performance in novel view synthesis by representing scenes with mixtures of Gaussians and utilizing differentiable rasterization.
-However, it typically requires large storage capacity and high VRAM, demanding the design of effective pruning and compression techniques.
-Existing methods, while effective in some scenarios, struggle with scalability and fail to adapt models based on critical factors such as computing capabilities or bandwidth, requiring to re-train the model under different configurations.
-In this work, we propose a novel, model-agnostic technique that organizes Gaussians into several hierarchical layers, enabling progressive Level of Detail (LoD) strategy.
-This method, combined with recent approach of compression of 3DGS, allows a single model to instantly scale across several compression ratios, with minimal to none impact to quality compared to a single non-scalable model and without requiring re-training.
+However, the high storage and VRAM requirements of 3DGS limit its deployment in resource-constrained environments, necessitating effective pruning and compression strategies to enable broader applicability.
+While existing methods are effective in certain scenarios, they lack scalability and adaptability to varying computational constraints, such as hardware capabilities or bandwidth limitations, often necessitating costly model retraining under different configurations.
+In this work, we introduce a novel and model-agnostic scalable compression framework for 3D Gaussian Splatting models.
+By leveraging our gradient-informed masking and compression algorithm, we construct a multi-level Gaussian hierarchy that efficiently balances quality and storage efficiency.
+At inference time, our method enables dynamic scaling across multiple compression levels without retraining, achieving comparable or superior results to state-of-the-art non-scalable compression approaches.
 We validate our approach on typical datasets and benchmarks, showcasing low distortion and substantial gains in terms of scalability and adaptability.
@@
 1 Introduction
+Figure 1: Our method builds a progressive Gaussian hierarchy, enabling adaptive LoD.
+Combined with compression, it supports multiple rates (eight in the figure), from high to low compression.
+The model can dynamically adjust based on criteria such as a target minimum FPS.
 Neural Radiance Fields (NeRFs) [26] have consistently boosted novel view synthesis, achieving state-of-the-art quality in scene reconstruction.
@@
 By leveraging differentiable rasterization and spherical harmonics for handling view-dependent effects, 3DGS achieves real-time performance even in complex real-world scenarios and at high-definition resolution.
-Figure 1: Our method constructs a progressive hierarchy of Gaussians, enabling progressive LoD.
-Combined with compression techniques, this allows the model to operate across several (eight in the figure) compression rates, ranging from extremely high to low.
-The model can dynamically adjust the rate based on criteria such as a target minimum FPS.
 3DGS offers competitive training times and reconstruction quality that can rival or even surpass that of NeRF-based models.
@@
 In fact, by adding or removing enhancement layers, we can easily adjust the trade-off between quality and performance.
-When combined with compression technique schemes, this approach enables scalable compression.
+When combined with a compression technique scheme, this approach enables scalable compression.
 Each level of detail can be seen as a distinct compression level, where higher LoDs corresponds to lower compression rate and lower LoDs corresponds to a higher compression rate.
@@
 This methods is general and compatible with popular 3DGS baselines.
-A fine-tuning phase ensures coherence and smoothness across the levels, while a final compression stage further reduce model size.
+A quantization-aware fine-tuning phase ensures coherence and smoothness across the levels, while a final compression step further reduce model size.
 In summary, our main contributions are:
   • A general method for progressive LoD (Sec.
-3.1): Leveraging a novel gradient-based iterative masking technique, we show the applicability of this method to a variety of popular 3DGS model (Sec.
-4.2) and how this approach effectively organizes primitives into a hierarchy of progressive levels with respect to other approaches (Sec.
+3.1): We propose a novel gradient-based iterative masking technique, which effectively organizes primitives into a hierarchy of progressive levels.
+Our approach demonstrates high scalability (Sec.
+4.4.1) and outperforms other methods in organizing the hierarchy (Sec.
 4.4.2).
-  • Smooth transition between levels (Sec.
-3.2): Our fine-tuning procedure offers a simple and effective way to reduce artifacts between levels, ensuring smooth transitions in visual fidelity across multiple LoDs (Sec.
-4.3) and boosting LoD quality (Sect.
+  • Smooth and continuous transitions between levels (Sec.
+3.2): Our fine-tuning procedure provides an effective method for achieving smoother transitions, significantly enhancing reconstruction quality by finding the optimal balance between different levels of detail.
+Transitions are further smoothed through a linear opacity interpolation mechanism, minimizing pop-up effects (Sec.
+4.3).
+  • Scalable Compression (Sec.
+3.3): By incorporating quantization-aware fine-tuning and entropy coding, our method achieves state-of-the-art compression-quality trade-offs (Sec.
+4.2), enabling navigation along the rate-distortion curve.
+This results in a negligible quality loss compared to non-compressed models (Sec.
 4.4.3).
-  • Scalable Compression (Sec.
-3.3): Our method integrates and extends state-of-the-art compression techniques, enabling scalable compression and adaptive rendering, with comparable and sometimes even superior performance of non-scalable methods (Sec.
-4.4.4).
-We extensively evaluate our approach on standard benchmark and compared it against the most recent state-of-the-art methods, showing comparable and even superior performances than non-scalable method.
-Our method allows models to scale across a wide range of compression levels, from low to very high, achieving up to 99.76% reduction in model size and 98.36% in the number of primitives in the highest compression rate with respect to non-compressed models, while preserving comparable full-size model quality at the lowest compression rate.
+We evaluate our method on various 3DGS architectures and compare it against the most recent state-of-the-art methods, showing comparable and even superior performances than non-scalable method.
+Our method allows models to scale across a wide range of compression levels, from low to very high, achieving up to 99.37% reduction in model size and 98.36% in the number of primitives in the highest compression rate with respect to non-compressed models, while preserving comparable full-size model quality at the lowest compression rate.
 2 Background and Related Work
@@
 LoD is a well-established technique in computer graphics, used to optimize performance and improve visual quality.
-It comprises three main components: constructing a scene representation at various levels of detail, implementing a policy for selecting the appropriate level based on the view, and enabling seamless transitions between levels to avoid artifacts or noticeable stuttering.
-Recent works have integrated LoD techniques with 3DGS, mainly to reduce aliasing artifacts [38] and improve performance, especially for large-scale datasets [23].
-These methods typically involve learning multiple levels of detail within the scene representation, followed by selective rendering according to heuristics like pixel coverage (i.e., the number of pixels occupied by a Gaussian projection on the image plane) and distance.
-Other work propose a hierarchical representation, often structured as a tree, as proposed by [24, 18].
+Recent work has extended LoD to 3DGS, primarily to mitigate aliasing artifacts [41] and improve performance, especially for large-scale datasets [24].
+These methods typically involve learning multiple levels of detail within the scene representation, followed by selective rendering according to heuristics such as pixel coverage (i.e., the number of pixels occupied by a Gaussian projection on the image plane) and distance from the camera.
+Other works propose a hierarchical representation, often structured as a tree, as proposed by [34, 19].
 This method dynamically implements LoD by selectively rendering Gaussians up to a specific level within the structure based on certain heuristics.
-In these approaches, LoD is typically defined at render time (dynamic LoD), which complicates its integration into compression contexts like ours, where a predefined, discrete set of progressive enhancement layers is essential.
-Additionally, these dynamic LoD methods often introduce significant overhead due to the need for tree traversal algorithms to construct levels of detail, as opposed to the simpler and faster operation of layer concatenation, which is more suitable for adaptive rendering in bandwidth-limited scenarios.
+Our technique is orthogonal to these works and can be integrated into dynamic LoD systems, where levels serve as an upper bound on active primitives.
+A related resolution-based LoD method [35] iteratively refines from a minimal representation, but suffers from three key limitations: high iteration count (30k per level), dependence on stored opacity values from previous levels, and rapid Gaussian proliferation at higher levels.
+In contrast, our approach supports an arbitrary number of levels without additional overhead, requiring only 30k iterations in total while maintaining a compact parameter count.
+Moreover, it generalizes across models and integrates compression for further efficiency.
 2.3 3DGS Compression
@@
 Current compression methods aim at achieving the highest quality at minimal storage size.
-Typically, compact representations involve post-hoc Gaussian pruning followed by compression stages, including quantization and entropy coding [1, 10, 30, 36].
-Some methods operate directly during training, aiming to limit the number of Gaussians via learned masks and optimized encoding schemes [14, 6, 22, 31, 34].
+Typically, compact representations involve post hoc Gaussian pruning followed by compression stages, including quantization (both during and after training) and entropy coding [2, 1, 32, 39].
+Some methods operate directly during training, with the aim of limiting the number of Gaussians through learned masks and optimized encoding schemes [15, 7, 23, 33, 37].
 Other studies enhance the representation by employing neural networks for on-the-fly Gaussian generation [25], tri-planes [38], or by using auxiliary structures such as hash grids to store spherical harmonics [7, 22], further reducing storage demands.
@@
 These anchor points, which are significantly fewer in number, are associated with a set of features, offsets, and scalings.
-A set of shallow neural networks process each anchor point and generates the parameters of view-adaptive Gaussians.
-However, it is important to note that none of these methods are scalable in practice.
-While different compression rates can be achieved, they require retraining the model from scratch or performing additional fine-tuning steps for each compression level.
-This limitation makes it difficult to adapt these model across a broad range of context, such as low bandwidth or varying hardware capabilities.
-In contrast, our method incorporates a scalable compression scheme: with a single training we can instantly obtain different compression levels, effectively navigating the rate-distortion curve.
-Furthermore, our method matches or even surpass dedicated, non-scalable compression techniques.
+A set of shallow neural networks processes each anchor point and generates the parameters of view-adaptive Gaussians.
+However, while these methods enable compression, they lack practical scalability.
+Adjusting compression rates requires either retraining from scratch or additional fine-tuning for each level, limiting their adaptability to scenarios with varying hardware constraints or bandwidth limitations.
+In contrast, our method incorporates a scalable compression scheme, enabling multiple compression levels from a single training run.
+This allows instant navigation of the rate-distortion curve without additional optimization.
+Moreover, our approach achieves performance comparable to, or even exceeding, dedicated non-scalable compression techniques.
 3 Gaussians on demand
@@
 1) Starting from a pre-trained model, we build a progressive hierarchy of Gaussians through iterative gradient-informed masking.
-2) A fine-tuning stage further enhances the model’s quality and reduces its memory footprint by culling unnecessary SH coefficients.
-3) The model is further compressed through codebook quantization.
-This section presents our proposed efficient and adaptive 3D rendering method using hierarchical Gaussian representations.
-The proposed approach extends the standard 3D Gaussian Splatting (3DGS) framework by introducing a hierarchical representation of the scene at multiple levels of detail.
-The objective is to develop a single, scalable model that can dynamically adapt to various criteria, such as computational resources or available bandwidth.
-We present a novel algorithm that implicitly constructs this hierarchy, enabling a single 3DGS model to learn multiple levels of detail.
-We first describe how to build a progressive hierarchy of Gaussians from a pre-trained 3DGS model (Sec.
-3.1) .
-Then, we show how to train and fine-tune Gaussians in different levels of detail (Sec.
+2) A quantization-aware fine tuning stage enhances the model’s quality and reduces its memory footprint.
+3) The model is further compressed through entropy coding.
+This section presents our framework for progressive LoD and scalable compression.
+Our approach extends the standard 3DGS framework by introducing a progressive hierarchy of Gaussians, with varying levels of detail.
+The objective is to develop a single, scalable model that can dynamically adapt to different compression levels, rendering budgets, and hardware or network constraints.
+We present a novel algorithm that constructs this hierarchy implicitly, enabling the model to learn multiple levels of detail.
+We first describe how to build the progressive hierarchy of Gaussians from a pre-trained 3DGS model (Sec.3.1), which forms the basis for constructing adaptable LoDs.
+Then, we explain how to train and fine-tune Gaussians for each level of detail (Sec.
 3.2).
-Finally, we describe our scalable compression and adaptive rendering steps (Sec.
+Finally, we describe our scalable compression method (Sec.
 3.3).
@@
 To implement progressive LoD, our objective is to organize the Gaussians of model $\mathcal{G}$ as a sequence of progressive levels $l\in\{0,\dots,L-1\}$ defined as
-$\mathcal{G}=(\mathcal{G}_{0},~{}\mathcal{E}_{1},\dots,\mathcal{E}_{l},\dots,\mathcal{E}_{L-1}),$ (1)
+$\mathcal{G}=(\mathcal{G}_{0},\leavevmode\nobreak\ \mathcal{E}_{1},\dots,\mathcal{E}_{l},\dots,\mathcal{E}_{L-1}),$ (1)
 where $\mathcal{G}_{0}$ denotes a set of starting Gaussians, and $\mathcal{E}_{l}$ is the set of enhancement Gaussians at level $l$.
@@
 The sets are organized progressively, ensuring that
-$\mathcal{G}_{0}\cap\mathcal{E}_{1},\dots,\cap~{}\mathcal{E}_{l},\dots,\cap~{}\mathcal{E}_{L-1}=\emptyset.$ (3)
+$\mathcal{G}_{0}\cap\mathcal{E}_{1},\dots,\cap\leavevmode\nobreak\ \mathcal{E}_{l},\dots,\cap\leavevmode\nobreak\ \mathcal{E}_{L-1}=\emptyset.$ (3)
 This guarantees that each level of detail progressively introduces new Gaussians, with no overlap among the layers.
@@
 We choose this progression since the rate-distortion functions generally follow a logarithmic law (see the theoretical rate-distortion curve for a Gaussian source [8]).
+This logarithmic progression is chosen to reflect the diminishing returns of adding additional Gaussians, but alternative progression schemes could be employed based on specific application constraints (e.g., computational capabilities of the rendering devices).
 To organize the Gaussians of the model into a hierarchical structure, we employ an iterative gradient-informed masking process.
 This approach follows a top-down strategy, starting with the construction of the higher enhancement levels first, namely, progressively building the model’s hierarchy from layer $l=L-1$ to layer $l=1$.
-The gradient of each parameter in the Gaussians serves as a critical and general indicator of its relevance to the overall model.
+The base layer $\mathcal{G}_{0}$ will consist of the remaining Gaussians at the end of the process.
+The gradient of each Gaussian parameter serves as a critical indicator of its relevance to the overall model.
 A Gaussian with a low gradient indicates minimal contribution to the loss function.
 Conversely, a high gradient suggests that the Gaussian is essential and has a significant impact on the loss.
-We can then leverage gradient information to build a hierarchy.
+We leverage this gradient information to construct the hierarchy.
 Specifically, to create the enhancement set $\mathcal{E}_{l}$, we accumulate gradients on $\mathcal{G}_{l}$ using the training set in order to select the $k_{l}$ Gaussians with the lowest importance score defined as the sum of the norm of the gradient of all parameters of each Gaussian:
 $\mathcal{E}_{l}=\texttt{lowest}_{k_{l}}\left\{\left\|\frac{\partial\mathcal{L}}{\partial\mathbf{\theta}_{i}}\right\|_{2}\right\}_{i\in\mathcal{G}_{l}},$ (6)
-where, with some abuse of notation, $\texttt{lowest}_{k_{l}}$ denotes the set of the $k_{l}$ Gaussians with the lowest importance.
+where $\texttt{lowest}_{k_{l}}$ selects the set of the $k_{l}$ Gaussians with the lowest gradient norm.
 Note that the difference $|\mathcal{G}_{l}|-|\mathcal{G}_{l-1}|$ is our $k_{l}$.
-The selected Gaussians will be associated with $\mathcal{E}_{l}$ and excluded (masked) from the next iterations.
+The selected Gaussians are associated with $\mathcal{E}_{l}$ and masked from further consideration in the next iterations.
 We reset the gradients and repeat the process.
-At the end of this iterative process, important Gaussians (with high gradient values) are placed at lower layers, while less important Gaussians are assigned to higher levels, where they model finer details and geometries.
 Although this phase requires $L-1$ steps, it tends to be very fast (1-2 minutes in total to build the hierarchy) as it only requires a single pass through the training set per step.
-3.2 Fine-Tuning and SHs culling
-This phase involves $F$ fine-tuning iterations designed to improve reconstruction quality and enforce a smooth transition between levels of detail.
+By the end of this process, important Gaussians (those with high gradients) are positioned at lower levels, contributing to coarse details, while less important Gaussians are placed at higher levels, modeling finer geometries and details.
+3.2 Quantization-Aware Fine-tuning
+This phase involves $F$ fine-tuning iterations.
 A random level $l$ is uniformly sampled at each iteration and a fine-tuning iteration is performed only on the $\mathcal{G}_{l}$ Gaussians associated with the selected level $l$.
-This optimization further improves reconstruction quality, especially at the lower levels.
-During fine-tuning, Gaussian parameters are reduced by selectively culling the spherical harmonics.
-Following the approach proposed by [31], SH coefficients are analyzed halfway through the fine-tuning phase to identify which primitives can be represented with 0, 1, 2, or 3 bands.
-The SH function is evaluated across all input views to determine whether a primitive requires view-dependent effects, based on the extent of color variation between views.
-If a primitive maintains a consistent color across all views, view-dependent effects are considered unnecessary.
-In addition, higher-order bands can be discarded if the color remains stable with lower-order SH coefficients.
-This phase significantly reduces the model’s memory cost (with negligible quality loss) since each Gaussian is represented by 59 32-bit floating point numbers, of which 45 are dedicated to the SH coefficients.
-3.3 Scalable Compression
-After fine-tuning, the compression step is performed.
-Compression is applied independently to the first LoD $\mathcal{G}_{0}$ and the following enhancements $\mathcal{E}_{l}$.
-Initially, a codebook is generated for each parameter (excluding the position of the primitives) using $K$-means clustering.
-In particular, following [31], codebooks of 256 entries are created for the opacity, the three scaling components, the real and imaginary parts of the rotation quaternion, the base color coefficients, and one for the 3-channel color component for each of the 15 SH coefficient groups.
-Codebook entries and positions are further quantized to 16-bit half-precision floats: in this way, we have achieved scalability since every LoD can be progressively decoded starting from the base one.
-It is worth pointing out that we move along the rate/distortion curve by just concatenating further enhancements.
-As shown in Sec.
-4, our results are very competitive with the recent state-of-the-art, with the added value of a single scalable model obtained with a single training process.
-In conclusion, GoDe achieves both scalable compression and progressive LoD.
-Scalable compression can be exploited in many application scenarios beyond the simple memory-saving result, e.g.
-adaptive HTTP streaming [32].
-The hierarchical organization of GoDe enhancements can be exploited for adaptive rendering at different LoD: specifically, the enhancements can be progressively splatted until a given target criterion, e.g.
-minimum frame rendering rate, is met.
-If rendering speed is the only goal, the GoDe compression phase can be skipped.
+This optimization enhances reconstruction quality, particularly at the lower levels, while also ensuring smoother transitions between them.
+During fine-tuning, we integrate quantization-aware training by quantizing spherical harmonic coefficients, opacity, scaling, and rotation parameters.
+Quantization is applied before the forward pass.
+To enable gradient flow through quantized parameters, a straight-through estimator is used, which allows backpropagation to proceed as if quantization were absent.
+More details are given in the Supplementary Sec.
+7.
+Since spherical harmonic coefficients are the most storage-intensive parameters, we further apply a L1 loss to encourage sparsity.
+Specifically, given the set of spherical harmonics coefficients $\mathcal{C}_{l}$ associated with the Gaussians in level $\mathcal{G}_{l}$ we introduce the penalty term:
+$\mathcal{L}_{\text{SH}}=\lambda\sum_{i}\|\mathbf{C}_{l,i}\|_{1}.$ (7)
+3.3 Scalable Compression and Adaptive Rendering
+After fine-tuning, a further compression step is applied.
+Compression is performed on the first LoD $\mathcal{G}_{0}$ as well as on the subsequent enhancement levels $\mathcal{E}_{l}$.
+This inherently enables scalable compression, where each level of detail can be interpreted as a compression stage.
+Specifically, lower LoDs correspond to higher compression rates, whereas higher LoDs yield lower compression.
+Each level can be progressively decoded starting from the previous one.
+Thus, the rate-distortion trade-off can be controlled simply by concatenating enhancement layers.
+As demonstrated in Sec.
+4, our approach achieves highly competitive results compared to the state-of-the-art, with the added benefit of relying on a single scalable model derived from a unified training process.
+This hierarchical organization can also be leveraged for adaptive rendering at varying levels of detail.
+More specifically, enhancement layers can be progressively splatted until a target criterion, such as a minimum frame rendering time threshold, is met.
+To prevent pop-up artifacts and ensure smooth transitions between LoDs, we introduce continuous transitions by linearly interpolating the opacity values of Gaussians in the subsequent enhancement layer.
+We define a vector of $N$ increasing interpolation factors:
+$\mathbf{t}=\{t_{1},t_{2},\dots,t_{N}\},\quad t_{i}\in[0,1],\quad t_{1}<t_{2}<\dots<t_{N}$ (8)
+and scale the Gaussians’ opacities in the enhancement layer accordingly.
+In conclusion, GoDe enables 3D Gaussian Splatting models to achieve scalable compression and adaptive rendering, providing state-of-the-art rate-distortion trade-offs.
 4 Experiments
@@
 The code was developed in Python 3.9 using PyTorch 2.1 and executed on a single NVIDIA A40 GPU.
-We set $L=8$ for all the configurations, $c_{\text{min}}=100,000$.
+We set $L=8$ for all the configurations, $c_{\text{min}}=100,000$ and $\lambda=10^{-2}$.
 The fine-tuning phase comprises $F=30,000$ iterations.
 All the other parameters remain unchanged from the original implementation.
-Our method was tested with three popular 3D Gaussian Splatting models: 3DGS [17], MCMC-3DGS [19], a variant of 3DGS that treats 3D Gaussians as samples from a Markov Chain Monte Carlo process, generally yielding slightly better results than 3DGS, and the neural-based Scaffold-GS [24].
-For Scaffold-GS, due to its unique architecture, we adopted a slightly different compression scheme.
-This approach is similar to our proposed GoDe method but excludes SHs culling since SHs are not present in the model.
-Instead, it incorporates iterative masking with quantization-aware fine-tuning and entropy coding.
-The iterative masking process follows a similar approach; however, rather than masking Gaussians, we mask anchor points directly, evaluating the gradients of their respective features.
-Quantization is performed during the fine-tuning phase: features are quantized to 8-bit unsigned integers, while the scaling, offset, and positional attributes are quantized to half-precision floating points similarly to [30] After fine-tuning, the model parameters are further compressed using ZIP, which combines LZ77 compression with Huffman entropy coding to achieve efficient and lossless storage reduction.
+Our method was evaluated with four popular 3D Gaussian Splatting models: 3DGS [18], MCMC-3DGS [20], a variant of 3DGS that treats 3D Gaussians as samples from a Markov Chain Monte Carlo process, generally yielding slightly better results than 3DGS, the neural-based Scaffold-GS [25] and the neural and dynamic-LoD-based Octree-GS [34].
+The spherical harmonic coefficients are quantized to 8-bit integers, while opacity, scaling, and rotation parameters are quantized to 16-bit floating-point values.
+Position parameters remain unquantized.
+For Scaffold-GS and Octree-GS the implementation is similar to the explicit models.
+Specifically, we quantize the anchor point features to 8-bit integers, while all the other parameters are quantized to 16-bit floating point values.
+To build the hierarchy, we use the gradient accumulated at each anchor point, constructing a progressive hierarchy of anchor points.
+After fine-tuning the model parameters are compressed using the LZMA algorithm.
 Table 1: Comparison of our methods against recent methods.
 All methods have been retrained and tested on the same hardware.
-For clarity, we present four compression levels out of 8 for our configurations, selected to match the compression rate of the other methods.
-These levels, labeled as Ours, are derived from the same model without the need for retraining (in contrast to the other approaches).
-Red, orange and yellow respectively indicates the first, second and third best result.
-Method Mip-NeRF360 Tanks&Temples Deep Blending
-PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size[MB] $\downarrow$ W $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size[MB] $\downarrow$ W $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size[MB] $\downarrow$ W $\downarrow$ FPS $\uparrow$
-HAC (low)  [6] 27.78 0.812 0.277 22.0 502 68 24.37 0.853 0.215 11.1 307 88 30.34 0.909 0.329 6.3 190 166
-HAC (high)  [6] 27.53 0.807 0.290 15.2 426 72 24.05 0.846 0.222 7.9 300 85 29.98 0.904 0.340 4.1 188 173
+For clarity, we present three compression levels out of 8 for our configurations, selected to match the compression rate of the other methods.
+These levels, labeled as Ours, are derived from the same model without the need for retraining.
+Red and yellow respectively indicates the first and the second best result.
+Mip-NeRF360 Tanks&Temples Deep Blending
+[…]
+27.43 0.811 0.271 49.0 1546 197 23.75 0.84 0.230 26.6 935 230 29.77 0.905 0.325 28.9 962 327
+27.30 0.804 0.293 20.6 603 282 23.65 0.834 0.244 13.6 437 337 29.78 0.904 0.333 14.8 451 465
+Ours 3DGS 26.63 0.776 0.341 8.6 242 391 23.12 0.813 0.278 7.0 208 470 29.70 0.900 0.348 7.4 212 622
+27.55 0.817 0.261 47.3 1546 143 24.04 0.845 0.217 26.6 938 178 29.73 0.903 0.326 26.7 965 267
+27.37 0.806 0.288 19.9 603 214 23.85 0.836 0.238 13.3 438 246 29.70 0.902 0.335 13.6 452 356
+Ours 3DGS-MCMC 26.75 0.775 0.338 8.4 242 299 23.22 0.812 0.278 6.8 208 359 29.59 0.898 0.350 7.0 212 459
+27.77 0.810 0.284 19.6 248 84 24.01 0.848 0.224 11.0 141 116 30.26 0.906 0.339 6.2 80 168
+27.61 0.803 0.297 14.1 179 94 23.96 0.844 0.231 8.9 128 128 30.02 0.901 0.350 4.5 58 180
+Ours Scaffold-GS 27.24 0.789 0.318 10.2 129 108 23.86 0.839 0.242 7.3 93 134 29.76 0.897 0.359 3.9 50 184
+27.73 0.813 0.271 19.5 262 71 24.40 0.856 0.242 13.2 173 65 30.10 0.902 0.338 7.3 103 135
+27.49 0.802 0.291 13.9 186 81 24.25 0.847 0.262 9.7 127 72 29.98 0.900 0.348 4.5 64 146
+Ours Octree-GS 27.01 0.780 0.322 10.0 133 92 24.01 0.834 0.286 7.1 93 78 29.87 0.898 0.352 4.0 56 146
+HAC (low) [7] 27.78 0.812 0.277 22.0 501 68 24.37 0.853 0.215 11.1 307 88 30.34 0.909 0.329 6.3 196 166
+HAC (high) [7] 27.17 0.798 0.306 11.5 371 74 24.03 0.843 0.237 7.4 261 82 29.92 0.903 0.350 3.9 180 166
 RDO [37] 27.05 0.801 0.288 23.3 1860 178 23.32 0.839 0.232 11.9 912 257 29.72 0.906 0.318 18.0 1478 200
-LightGaussian  [10] 27.24 0.810 0.273 51.0 2197 161 23.55 0.839 0.235 28.5 1210 225 29.41 0.904 0.329 43.2 956 348
+Reduced-3DGS [33] 27.19 0.810 0.267 29.3 1436 234 23.55 0.843 0.223 14.0 656 358 29.70 0.907 0.315 18.3 990 304
+Comp-GS [23] 27.28 0.802 0.283 16.4 485 165 23.62 0.837 0.248 9.8 241 286 29.89 0.904 0.336 10.0 246 254
 SOG [28] 27.08 0.799 0.277 40.3 2176 132 23.56 0.837 0.221 22.8 1242 227 29.26 0.894 0.336 17.7 890 236
+EAGLES [15] 27.13 0.809 0.278 57.1 1290 156 23.20 0.837 0.241 28.9 651 227 29.75 0.910 0.318 52.4 1192 144
+LightGuassian [11] 27.24 0.810 0.273 51.0 2197 161 23.55 0.839 0.235 28.5 1211 225 29.41 0.904 0.329 43.2 956 348
 Compact-3DGS [22] 27.02 0.800 0.287 29.1 1429 108 23.41 0.836 0.238 20.9 842 147 29.76 0.905 0.324 23.8 1053 144
-Reduced-3DGS [31] 27.19 0.810 0.267 29.3 1436 234 23.55 0.843 0.223 14.0 656 358 29.70 0.907 0.315 18.3 990 304
-EAGLES [14] 27.13 0.809 0.278 57.1 1290 156 23.20 0.837 0.241 28.9 651 227 29.75 0.910 0.318 52.4 1192 144
-Comp-GS [22] 27.28 0.802 0.283 16.4 485 165 23.62 0.837 0.248 9.8 241 286 29.89 0.904 0.336 10.0 246 253
-27.75 0.810 0.284 19.7 453 98 24.01 0.846 0.226 11.0 266 142 30.34 0.909 0.331 8.4 63 212
-27.59 0.803 0.297 14.1 369 107 23.95 0.843 0.234 8.9 230 156 30.29 0.908 0.334 7.2 56 208
-27.34 0.793 0.316 10.3 288 127 23.85 0.837 0.244 7.3 196 176 30.14 0.905 0.343 5.3 44 225
-Ours w/ Scaffold-GS 27.07 0.780 0.336 7.5 221 146 23.72 0.830 0.258 5.9 165 189 29.76 0.897 0.358 3.9 33 236
-27.42 0.815 0.263 43.9 1546 146 23.97 0.842 0.220 22.2 938 216 29.71 0.901 0.323 19.5 931 263
-27.23 0.804 0.289 20.0 603 222 23.76 0.831 0.241 11.8 438 311 29.70 0.901 0.326 11.0 491 345
-26.99 0.790 0.312 13.4 381 268 23.49 0.821 0.259 8.6 301 366 29.66 0.901 0.331 8.3 357 391
-Ours w/ 3DGS-MCMC 26.59 0.772 0.339 9.0 242 317 23.06 0.806 0.281 6.3 208 426 29.56 0.897 0.348 4.8 189 488
-27.27 0.807 0.273 40.5 1546 189.1 23.76 0.839 0.231 21.1 938 269 29.73 0.904 0.327 19.8 931 302
-27.16 0.801 0.295 18.3 603 270 23.66 0.832 0.245 11.6 438 383 29.73 0.903 0.334 10.1 491 425
-26.93 0.791 0.315 12.0 381 317 23.48 0.824 0.259 8.6 301 439 29.74 0.902 0.340 7.2 357 492
-Ours w/ 3DGS 26.55 0.773 0.341 7.9 242 371 23.13 0.812 0.279 6.4 208 509 29.67 0.899 0.348 5.1 189 564
 4.2 Quantitative Results
-Figure 3: Rate-distortion curves (left), PSNR/W (center, where W is the number of primitives) and PSNR/FPS (right) curve on Mip-NeRF360.
+Figure 3: Rate/distortion (above) and PSNR/FPS (below) curves on the three main datasets.
 Our approach achieves competitive results with both neural-based Gaussian models and explicit ones while maintaining significantly higher speed at comparable quality.
@@
 We divide our analysis of the results between neural-based 3DGS models and explicit ones, as the former generally provides superior rate-distortion performance.
+Notably, the models obtained with our method are highly scalable, ranging from an extremely high compression rate of a few megabytes to an extremely low one.
 Neural-based 3DGS
@@
 Scaffold-GS provides slightly higher reconstruction quality than 3DGS (approximately +0.3 dB) while reducing storage by about 5$\times$.
-By applying our method to Scaffold-GS we observe that it delivers comparable results, slightly outperforming HAC on the MipNeRF-360 dataset, while HAC remains superior on the Tanks&Temples dataset.
+By applying our method to Scaffold-GS and Octree-GS we observe that it delivers comparable results, slightly outperforming HAC on the Mip-NeRF360 and Tanks&Temples datasets, while HAC remains slightly superior only on the DeepBlending dataset.
 We consider our results competitive, as our scalable compression method does not require re-training, unlike other methods that need to train a separate model with different hyperparameters.
@@
 There is one point where Comp-GS [23] slightly outperforms our method, but this approach scales less effectively at higher compression rates.
-In this case, our approach also offers the advantage of utilizing a single scalable model, which eliminates the need for re-training when adapting to different compression rates.
-Notably, the models obtained with our method are highly scalable, ranging from an extremely high compression rate of a few megabytes to an extremely low one.
-4.3 Visualizing LoDs
-In Fig.
-4 we visualize some LoD levels for various scenes.
-For each LoD level, we report the average frame rate computed for rendering the image of the test set, the number of Gaussians, the average PSNR, and the model size.
-The transition between LoDs is smooth thanks to the fine-tuning process with random LoD sampling.
-Moreover, we observe that lower LODs are extremely fast and small, with rendering speeds exceeding 800 FPS and model sizes as low as 2.4MB for the drjohnson scene from the DeepBlending dataset (in comparison with the vanilla-3DGS which renders the same scene at 146FPS and requires 780MB of storage).
-Similar performances are achieved across other datasets.
-It is important to note that lower LoDs tend to lose fine details, particularly in high-frequency areas such as reflections, transparency handling, and background elements.
-This is particulary noticeable in the kitchen scene, where at lower levels of detail the glass bottle in the background is either not rendered at all or only partially and becomes progressively more defined at higher levels.
-Another example can be seen in the counter scene, with the reflection on the dishes.
-Other examples are the paining in the drjohnson scene and the trees in the background of the $truck$ scene.
-This is a significant result, demonstrating the effectiveness of the method, which automatically learns a hierarchy and preserves the most important details at each level, minimizing the loss of detail to secondary aspects.
-This is further illustrated in Fig.
-5, where we render only the Gaussian associated with an enhancement level.
-As the level increases, we observe that the Gaussians focus on more complex geometry and finer details.
+In this case as-well, our approach also offers the advantage of utilizing a single scalable model, which eliminates the need for re-training when adapting to different compression rates.
+4.3 Visualizing LoDs and Transitions
 Figure 4: Visualization of LoDs.
@@
 Starting from top to bottom we have: kitchen, counter, drjohnson and truck scenes.
-Figure 5: Visualization of enhancement levels. We render only the Gaussians belonging to the specified enhancement level.
+In Fig.
+4 we visualize some LoD levels for various scenes.
+For each LoD level, we report the average frame rate computed for rendering the image of the test set, the number of Gaussians, the average PSNR, and the model size.
+The transition between LoDs is smooth, due to the fine-tuning process with random LoD sampling.
+Moreover, we observe that lower LODs are extremely fast and small, with rendering speeds exceeding 900 FPS and model sizes as low as 3.7 MB for the drjohnson scene from the DeepBlending dataset (compared to the vanilla-3DGS which renders the same scene at 146FPS and requires 780MB of storage).
+Similar performance is achieved in other datasets.
+It is important to note that lower LoDs tend to lose fine details, particularly in high-frequency areas such as reflections, transparency handling, and background elements.
+This is particularly noticeable in the kitchen scene, where at lower levels of detail the glass bottle in the background is either not rendered at all or only partially and becomes progressively more defined at higher levels.
+Another example can be seen in the counter scene, with the reflection on the dishes.
+Other examples are the paining in the drjohnson scene and the trees in the background of the $truck$ scene.
+This is a significant result, demonstrating the effectiveness of the method, which automatically learns a hierarchy and preserves the most important details at each level, minimizing the loss of detail to secondary aspects.
+In Fig.
+5, we illustrate a seamless transition between two adjacent levels of detail, achieved by linearly interpolating the opacity values of the nearest enhancement layer Gaussians.
+This approach further minimizes pop-up artifacts.
+Additional visualizations, including videos, are provided in the supplementary material.
+Figure 5: Continuous transitions between two adjacent levels, with eight linear interpolating factors on Garden scene.
+Figure 6: Ablation on the number of levels on Mip-NeRF360.
+Figure 7: Comparison among pruning methods on Mip-NeRF 360.
+Figure 8: Comparison among our method with Scaffold-GS, a non-scalable variant of the same method and HAC.
 4.4 Ablations
@@
 Since the iterative masking process is very fast, we opted to use the sum of the gradient norms of the parameters as a score and perform the masking iteratively.
-Figure 6: Ablation on the number of levels on Mip-NeRF360.
-Figure 7: Comparison among pruning methods on Mip-NeRF 360.
-Figure 8: Comparison among our method with Scaffold-GS, a non-scalable variant of the same method and HAC. This constitutes an upper bound to our scalable method.
 Table 2: Effects of fine-tuning and compression on Mip-NeRF360.
-LoD W Mask Only Fine Tuning Fine-Tuning + Compression
+Mask Only Fine Tuning Fine-Tuning + Compression
 […]
-0 100 18.52 0.553 24.1 25.20 0.719 24.1 25.24 0.717 3.5
-1 155 19.90 0.606 36.7 26.06 0.752 36.7 26.03 0.749 5.3
-2 242 21.52 0.662 56.4 26.65 0.778 56.4 26.55 0.773 7.9
-3 381 23.28 0.715 89.1 26.84 0.787 89.1 26.93 0.791 12.0
-4 603 25.00 0.760 142.8 27.08 0.796 142.8 27.16 0.801 18.3
-5 962 26.45 0.793 216.6 27.33 0.807 69.8 27.25 0.806 27.5
-6 1546 27.23 0.810 365.9 27.45 0.813 227.2 27.27 0.807 40.5
-7 2497 27.49 0.815 562.2 27.47 0.814 365.6 27.29 0.808 57.9
-AVG 811 23.67 0.714 186.7 26.76 0.783 126.5 26.72 0.781 21.6
+0 100 18.52 0.553 24.1 25.20 0.719 24.1 25.28 0.719 3.8
+1 155 19.90 0.606 36.7 26.06 0.752 36.7 26.10 0.751 5.6
+2 242 21.52 0.662 56.4 26.65 0.778 56.4 26.63 0.776 8.6
+3 381 23.28 0.715 89.1 26.84 0.787 89.1 27.05 0.793 13.3
+4 603 25.00 0.760 142.8 27.08 0.796 142.8 27.30 0.804 20.6
+5 962 26.45 0.793 216.6 27.33 0.807 69.8 27.40 0.809 32.1
+6 1546 27.23 0.810 365.9 27.45 0.813 227.2 27.43 0.811 49.0
+7 2497 27.49 0.815 562.2 27.47 0.814 365.6 27.42 0.811 73.7
 Figure 9: Effects of fine-tuning on the train scene of Tanks&Temples dataset. In the first row, LoDs without fine-tuning; on the second, LoDs after fine-tuning.
@@
 2, the fine-tuning phase significantly improves reconstruction quality, particularly for lower LoDs, while slightly reducing quality in the two highest LoDs, of about 0.02 dB.
-Regarding compression, we observe a reduction in storage ranging from $99.37\%$ to $89.69\%$, depending on the level considered.
-The impact on reconstruction quality is minimal, with a loss of 0.03 dB to 0.18 dB across different levels.
+Regarding compression, we observe a reduction in storage ranging from $99.37\%$ to $86.88\%$, depending on the level considered.
+The impact on reconstruction quality is negligible, with a loss of 0.05 dB in the highest LoD.
 In Fig.
@@
 8).
-This procedure is repeated across 8 levels, down to a minimum threshold of $100,000$ parameters, with each level acting as an independent model.
+This procedure is repeated across 8 levels, down to a minimum threshold of $50,000$ parameters, with each level acting as an independent model.
 Interestingly, our method sets a new performance benchmark, outperforming HAC and underline the effectiveness of iterative, gradient-based pruning for this task.
@@
 This is due to the imposed hierarchy, which requires optimizing each level in relation to the others, thus introducing a trade-off between higher and lower levels.
-5 Limitation and Future Work
-Our approach demonstrates effectiveness in scalable compression tasks, matching or even surpassing the performance of non-scalable compression methods.
-However, it also inherits certain limitations typical of top-down pruning and compression approaches, such as the requirement for a pre-trained model with a well-established geometric prior.
-In future work, we intend to explore the potential for a bottom-up approach, integrating the hierarchical structure during the training of the Gaussian model.
 5 Conclusion
@@
 This framework also enables adaptive rendering: given the hierarchical-layered structure of the model, it is possible to dynamically select Gaussians up to a certain level to meet specific criteria, such as a target frame rate.
-Additionally, a notable advantage of our approach is its generality: relying on widely applicable techniques, such as gradient-informed pruning, our method can be integrated into popular 3DGS models.
+Additionally, a notable advantage of our approach is its generality: relying on widely applicable techniques, such as gradient-informed pruning and quantization-aware training, our method can be integrated into popular 3DGS models.
+Supplementary Material
+In Sec.
+6 we present the pseudocode for building the hierarchy and the fine-tune process.
+In Sec.
+7 we provide more details about the quantization-aware training.
+In Sec.
+8 we discuss alternative sampling strategy for the fine-tuning phase.
+In Sec.
+9 we provide a comparison of our method with and without compression enabled.
+In Sec.
+10 we present additional results, including a more comprehensive comparison with state-of-the-art methods (Sec.
+10.1), a complete visualization of levels across all scenes (Sec.
+10.2) and per-scenes results (Sec.
+10.3).
+6 Pseudocode
+Here, we present the code for constructing the hierarchical Gaussian model and the fine-tuning process.
+Following Eq.
+1, a Hierarchical Gaussian model $\mathcal{G}$ is defined as a base layer $\mathcal{G}_{0}$ and a list of enhancement layers $\mathcal{E}_{s}$.
+The process begins with the computation of a progression that determines the number of primitives at each level, as described in Alg.
+1.
+Given a minimum number of Gaussians, $c_{\text{min}}$, which defines the number of gaussians in the base layer, and the maximum number of Gaussians, $c_{\text{max}}$, which sets the total count, we compute the progression of $k_{s}$.
+In our experiments, we impose a limit on the total number of Gaussians, capping it at $0.75\%$ of the original model’s count.
+This progression is chosen based on the observation that rate-distortion curves typically follow a logarithmic pattern.
+However, it is important to note that our algorithm is independent of this assumption.
+We then proceed with constructing the hierarchy in a top-down manner, as outlined in Alg.
+2, followed by the final fine-tuning step, which is described in Alg.
+3.
+Algorithm 1 Compute $k_{s}$
+1:number of starting Gaussians $c_{\text{min}}$, maximum number of Gaussians $c_{\text{max}}$, number of levels $L$
+2:list of Gaussian counts per level $k_{s}$
+3:$b\leftarrow\exp\left(\frac{\log(c_{\text{max}})-\log(c_{\text{min}})}{L-1}\right)$
+4:$arr\leftarrow[0,1,\dots,L-1]$
+5:$progr\leftarrow\text{floor}((b^{arr})\times c_{\text{min}})$
+6:$k_{s}\leftarrow[c_{\text{min}}]$
+7:for $i\leftarrow 1$ to $L-1$ do
+8:       append($k_{s},progr[i]-progr[i-1]$)
+9:end for
+10:return $k_{s}$
+7 Implementation Details on QA Fine-Tuning
+To implement quantization, we used PyTorch’s quantization libraries.
+The Quantization-Aware Training in PyTorch employs symmetric quantization with fixed intervals and is based on FakeQuantization, which emulates quantization without losing the floating-point representation.
+This technique is general and easily adaptable to all 3DGS models.
+Specifically, for explicit models (3DGS and 3DGS-MCMC), we quantized the Spherical Harmonics coefficients to 8-bit integers, while the remaining parameters were quantized to 16-bit floating point, leaving the position values unchanged.
+For Scaffold-GS and Octree-GS models, the features of the anchor points were quantized to 8-bit integers, while all other parameters, including position, were kept in 16-bit floating point.
+In both cases, an auxiliary L1 loss was used: for explicit models, it was applied to the SH coefficients, and for neural models, it was applied to the anchor point features.
+During the forward pass, quantization involves rounding and clipping.
+Given a floating-point value $x$, the quantized value $x_{q}$ is computed as:
+$x_{q}=\text{clip}(\text{round}(\frac{x}{\text{scale}}+\text{zero\_point}),q_{\text{min}},q_{\text{max}})$
+where $q_{\text{min}}$ and $q_{\text{max}}$ are determined by the number of bits used, scale is defined as $\frac{x_{\text{max}}-x_{\text{min}}}{q_{\text{max}}-q_{\text{min}}}$, and zero_point is calculated as $q_{\text{min}}-\left(\frac{x_{\text{min}}}{\text{scale}}\right)$.
+After quantization, the value is dequantized as:
+[…]
+During the backward pass, a straight-through estimator is implemented to bypass the rounding operation.
+Algorithm 2 Iterative Masking
+1:Pre-trained Gaussian Model $\mathcal{G}_{p}$, number of levels $L$, progression $k_{s}$, training set $\mathcal{D}_{\text{train}}$
+2:Hierarchical Gaussian model $(\mathcal{G}_{0},\mathcal{E}_{s})$
+3:$\mathcal{E}_{s}\leftarrow$ empty list of size $L-1$
+4:$\mathcal{G}_{L-1}\leftarrow\mathcal{G}_{p}$
+5:for $l\leftarrow L-1$ to $1$ do
+6:       computeGradients($\mathcal{G}_{l}$, $\mathcal{D}_{\text{train}}$)
+7:       $k_{l}\leftarrow k_{s}[l]$
+8:       $\mathcal{E}_{l}\leftarrow\texttt{lowest}_{k_{l}}\left\{\left\|\frac{\partial\mathcal{L}}{\partial\mathbf{\theta}_{i}}\right\|_{2}\right\}_{i\in\mathcal{G}_{l}}$
+9:       append($\mathcal{E}_{l}$, $\mathcal{E}_{s}$)
+10:       $\mathcal{G}_{l-1}\leftarrow$ maskGaussians($\mathcal{E}_{l}$, $\mathcal{G}_{l}$)
+11:end for
+12:reverse($\mathcal{E}_{s}$)
+13:return $(\mathcal{G}_{0},\mathcal{E}_{s})$
+Algorithm 3 Fine Tune
+1:Hierarchical Gaussian Model $(\mathcal{G}_{0},\mathcal{E}_{s})$, training set $\mathcal{D}_{\text{train}}$, number of iterations $F$, boolean compress
+2:Fine-tuned Hierarchical Model ($\mathcal{G}_{0},\mathcal{E}_{s}$)
+3:for $i\leftarrow 0$ to $F$ do
+4:       Random sample viewpoint $v$ and level $l\in\{0,|\mathcal{E}_{s}|\}$
+5:       $\mathcal{G}_{l}\leftarrow\mathcal{G}_{0}\cup\bigcup_{i=1}^{l}\mathcal{E}_{l}$
+6:       render($\mathcal{G}_{l},v$)
+7:       compute loss and optimize $\mathcal{G}_{l}$
+8:       if compress then
+9:             compressModel($\mathcal{G}_{0},\mathcal{E}_{s}$)
+10:       end if
+11:end for
+12:return ($\mathcal{G}_{0},\mathcal{E}_{s}$)
+8 Random Sampling vs Weighted Sampling
+Table 3: Comparison among sampling strategies on Mip-NeRF360 and 3DGS.
+level W Uniform Sampling Weighted Sampling (G=5) Weighted Sampling (G=8) Weighted Sampling (G=10)
+[…]
+LoD 0 100 25.20 0.719 0.408 24.72 0.706 0.417 24.12 0.693 0.426 23.30 0.675 0.436
+LoD 1 155 26.06 0.752 0.372 25.70 0.742 0.379 25.26 0.732 0.386 24.70 0.721 0.392
+LoD 2 242 26.65 0.778 0.340 26.41 0.771 0.345 26.13 0.764 0.349 25.90 0.760 0.351
+LoD 3 381 27.08 0.796 0.312 26.94 0.792 0.315 26.89 0.791 0.315 26.77 0.788 0.316
+LoD 4 603 27.33 0.807 0.291 27.37 0.808 0.289 27.35 0.807 0.289 27.34 0.807 0.288
+LoD 5 962 27.45 0.813 0.276 27.56 0.815 0.272 27.58 0.815 0.271 27.58 0.815 0.270
+LoD 6 1546 27.47 0.814 0.269 27.58 0.816 0.264 27.61 0.817 0.263 27.63 0.817 0.261
+LoD 7 2497 27.47 0.815 0.267 27.57 0.816 0.262 27.60 0.817 0.260 27.62 0.817 0.259
+AVG 811 26.84 0.787 0.317 26.73 0.783 0.318 26.57 0.779 0.320 26.36 0.775 0.322
+In this section, we compare the random sampling approach with alternative sampling strategies.
+During the fine-tuning phase, the levels are sampled uniformly, following $l\sim\text{Uniform}\{0,L-1\}$, where $L$ denotes the number of levels.
+Additionally, we explored a weighted sampling strategy that assigns higher probabilities to upper levels.
+While this method improves the quality of the upper levels, it significantly degrades the quality of the lower ones, as shown in Tab.
+3.
+To implement the weighted sampling, we utilize Eq.
+5 and transform the progression of $k_{s}$ into probabilities, normalizing them between 0 and 1.
+This results in upper levels receiving higher weights, while lower levels are assigned lower weights, making them less probable.
+The degree of emphasis on upper levels is controlled by the denominator $(L-1)$ (here redefined as $G$ for clarity).
+Increasing $G$ accentuates the bias toward upper levels, whereas reducing tends to an uniform sampling scheme.
+We tested with G up to 10, observing that uniform sampling of levels generally produces better results, establishing it as an optimal trade-off between higher and lower levels.
+9 Compression vs Not Compression
+In Tab.
+456 we present the results for all 8 levels and models, both with and without compression.
+The observed quality loss due to compression is negligible compared to the substantial gains in storage efficiency, especially for explicit models.
+The maximum quality degradation occurs with MCMC-3DGS on the Mip-NeRF360 dataset, amounting to -0.19dB (LoD 8), while on Tanks&Temples, it is even less significant, reaching only -0.12dB (LoD 8).
+Remarkably, on the DeepBlending dataset, compression results in a general improvement in quality, with gains of up to +0.23dB (LoD 8), with the exception of Octree-GS.
+Table 4: Compressed vs Not Compressed models on Mip-NeRF360
+Method LoD W Compressed Not Compressed
+[…]
+Ours 3DGS LoD 0 100 25.28 0.719 0.407 3.8 25.20 0.719 0.408 23.7
+LoD 1 155 26.10 0.751 0.371 5.6 26.06 0.752 0.372 35.9
+LoD 2 242 26.63 0.776 0.341 8.6 26.65 0.778 0.340 56.3
+LoD 3 381 27.05 0.793 0.314 13.3 27.08 0.796 0.312 89.0
+LoD 4 603 27.30 0.804 0.293 20.6 27.33 0.807 0.291 141.5
+LoD 5 962 27.40 0.809 0.278 32.1 26.84 0.787 0.317 226.5
+LoD 6 1546 27.43 0.811 0.271 49.0 27.45 0.813 0.276 364.8
+LoD 7 2497 27.42 0.811 0.268 73.7 27.47 0.814 0.269 590.6
+Ours 3DGS-MCMC LoD 0 100 25.40 0.718 0.406 3.7 25.26 0.718 0.407 23.7
+LoD 1 155 26.22 0.750 0.370 5.6 26.15 0.752 0.369 35.9
+LoD 2 242 26.75 0.775 0.338 8.4 26.77 0.779 0.334 56.3
+LoD 3 381 27.12 0.794 0.310 12.9 27.23 0.800 0.305 89.0
+LoD 4 603 27.37 0.806 0.288 19.9 27.52 0.814 0.280 141.5
+LoD 5 962 27.50 0.814 0.271 30.8 27.69 0.823 0.262 226.5
+LoD 6 1546 27.55 0.817 0.261 47.3 27.74 0.826 0.252 364.8
+LoD 7 2497 27.55 0.819 0.257 72.2 27.74 0.827 0.248 590.6
+Ours Scaffold-GS LoD 0 50 24.97 0.710 0.409 4.0 25.03 0.711 0.408 15.1
+LoD 1 68 25.98 0.743 0.375 5.4 26.03 0.744 0.374 20.7
+LoD 2 94 26.68 0.768 0.345 7.4 26.72 0.769 0.344 28.6
+LoD 3 129 27.24 0.789 0.318 10.2 27.27 0.790 0.317 39.5
+LoD 4 179 27.61 0.803 0.297 14.1 27.63 0.804 0.296 54.3
+LoD 5 248 27.77 0.810 0.284 19.6 27.81 0.811 0.283 75.7
+LoD 6 346 27.84 0.812 0.278 27.2 27.88 0.813 0.277 104.9
+LoD 7 483 27.87 0.813 0.276 37.9 27.89 0.814 0.275 146.2
+Ours Octree-GS LoD 0 50 24.61 0.689 0.426 3.7 24.63 0.690 0.425 14.7
+LoD 1 69 25.55 0.722 0.391 5.2 25.57 0.723 0.390 20.2
+LoD 2 96 26.33 0.752 0.357 7.1 26.36 0.753 0.356 28.1
+LoD 3 133 27.01 0.780 0.322 10.0 27.05 0.781 0.321 39.1
+LoD 4 186 27.49 0.802 0.291 13.9 27.53 0.802 0.291 54.7
+LoD 5 262 27.73 0.813 0.271 19.5 27.78 0.814 0.270 77.0
+LoD 6 371 27.85 0.818 0.260 27.5 27.89 0.818 0.260 108.9
+LoD 7 527 27.87 0.818 0.257 39.0 27.91 0.819 0.257 154.7
+Table 5: Compressed vs Not Compressed models on Tanks&Temples.
+Method LoD W Compressed Not Compressed
+[…]
+Ours 3DGS LoD 0 100 21.82 0.770 0.330 3.6 21.84 0.775 0.323 23.7
+LoD 1 144 22.62 0.795 0.302 5.0 22.64 0.801 0.295 34.0
+LoD 2 208 23.12 0.813 0.278 7.0 23.17 0.819 0.271 49.1
+LoD 3 301 23.46 0.826 0.258 9.7 23.55 0.832 0.251 71.2
+LoD 4 438 23.65 0.834 0.244 13.6 23.75 0.841 0.236 103.6
+LoD 5 639 23.73 0.838 0.235 19.2 23.83 0.845 0.227 151.2
+LoD 6 936 23.75 0.840 0.230 26.6 23.87 0.847 0.222 221.5
+LoD 7 1376 23.75 0.841 0.227 36.8 23.87 0.848 0.219 325.5
+Ours 3DGS-MCMC LoD 0 100 21.92 0.769 0.333 3.5 21.74 0.771 0.328 23.7
+LoD 1 144 22.69 0.793 0.303 4.9 22.55 0.796 0.296 34.0
+LoD 2 208 23.22 0.812 0.278 6.8 23.14 0.816 0.269 49.1
+LoD 3 301 23.62 0.826 0.255 9.5 23.61 0.832 0.245 71.2
+LoD 4 438 23.85 0.836 0.238 13.3 23.89 0.844 0.226 103.6
+LoD 5 640 23.99 0.841 0.226 18.7 24.06 0.851 0.212 151.2
+LoD 6 938 24.04 0.845 0.217 26.6 24.11 0.854 0.203 221.5
+LoD 7 1380 24.05 0.845 0.215 37.4 24.12 0.856 0.197 325.5
+Ours Scaffold-GS LoD 0 50 23.13 0.806 0.295 3.9 23.13 0.806 0.295 15.1
+LoD 1 61 23.51 0.821 0.273 4.8 23.51 0.821 0.273 18.5
+LoD 2 75 23.72 0.832 0.256 5.9 23.74 0.832 0.256 22.8
+LoD 3 93 23.86 0.839 0.242 7.3 23.88 0.839 0.242 28.0
+LoD 4 114 23.96 0.844 0.231 8.9 23.98 0.844 0.231 34.5
+LoD 5 141 24.01 0.848 0.224 11.0 24.04 0.848 0.224 42.5
+LoD 6 174 24.04 0.850 0.219 13.5 24.07 0.850 0.220 53.7
+LoD 7 215 24.04 0.851 0.218 16.6 24.07 0.851 0.218 65.5
+Ours Octree-GS LoD 0 50 23.08 0.798 0.342 3.9 23.10 0.800 0.340 14.7
+LoD 1 68 23.68 0.819 0.312 5.3 23.70 0.820 0.310 20.0
+LoD 2 93 24.01 0.834 0.286 7.1 24.03 0.840 0.280 27.3
+LoD 3 127 24.25 0.847 0.262 9.7 24.26 0.850 0.260 37.2
+LoD 4 173 24.40 0.856 0.242 13.2 24.42 0.860 0.240 50.7
+LoD 5 235 24.50 0.863 0.227 17.9 24.52 0.860 0.230 69.2
+LoD 6 321 24.56 0.867 0.218 24.2 24.58 0.870 0.220 94.3
+LoD 7 438 24.56 0.868 0.215 32.8 24.59 0.870 0.210 128.6
+Table 6: Compressed vs Not Compressed models on DeepBlending.
+Method LoD W Compressed Not Compressed
+[…]
+Ours 3DGS LoD 0 100 29.22 0.888 0.372 3.7 29.01 0.888 0.373 23.7
+LoD 1 137 29.57 0.896 0.358 5.3 29.44 0.897 0.355 36.5
+LoD 2 189 29.70 0.900 0.348 7.4 29.60 0.902 0.343 56.3
+LoD 3 259 29.78 0.902 0.339 10.5 29.62 0.905 0.334 87.0
+LoD 4 357 29.78 0.904 0.333 14.8 29.60 0.905 0.328 134.5
+LoD 5 491 29.78 0.904 0.328 20.7 29.58 0.905 0.324 208.1
+LoD 6 676 29.77 0.905 0.325 28.9 29.55 0.905 0.322 322.0
+LoD 7 931 29.77 0.905 0.324 39.8 29.54 0.904 0.322 498.6
+Ours 3DGS-MCMC LoD 0 100 29.12 0.887 0.376 3.6 28.95 0.888 0.373 23.7
+LoD 1 137 29.44 0.894 0.361 5.0 29.41 0.896 0.356 36.5
+LoD 2 189 29.59 0.898 0.350 7.0 29.58 0.901 0.343 56.3
+LoD 3 259 29.65 0.900 0.342 9.8 29.63 0.903 0.333 87.0
+LoD 4 357 29.70 0.902 0.335 13.6 29.64 0.904 0.326 134.5
+LoD 5 491 29.72 0.902 0.329 19.1 29.62 0.904 0.321 208.1
+LoD 6 676 29.73 0.903 0.326 26.7 29.58 0.904 0.319 322.0
+LoD 7 931 29.71 0.902 0.325 37.5 29.55 0.903 0.318 498.6
+Ours Scaffold-GS LoD 0 50 29.76 0.897 0.359 3.9 29.73 0.896 0.360 15.1
+LoD 1 58 30.02 0.901 0.350 4.5 29.97 0.900 0.351 17.6
+LoD 2 68 30.16 0.904 0.344 5.3 30.11 0.903 0.345 20.5
+LoD 3 80 30.26 0.906 0.339 6.2 30.20 0.905 0.340 24.0
+LoD 4 93 30.31 0.908 0.335 7.2 30.27 0.907 0.336 28.2
+LoD 5 109 30.36 0.909 0.332 8.4 30.31 0.907 0.333 34.0
+LoD 6 127 30.38 0.909 0.330 9.8 30.33 0.908 0.331 38.7
+LoD 7 149 30.38 0.909 0.329 11.4 30.31 0.908 0.330 46.0
+Ours Octree-GS LoD 0 50 29.67 0.895 0.358 3.6 29.72 0.895 0.358 14.7
+LoD 1 56 29.87 0.898 0.352 4.0 29.92 0.898 0.351 16.6
+LoD 2 64 29.98 0.900 0.348 4.5 30.04 0.900 0.347 18.7
+LoD 3 72 30.04 0.901 0.344 5.1 30.10 0.901 0.343 21.1
+LoD 4 81 30.07 0.902 0.341 5.7 30.13 0.902 0.341 23.7
+LoD 5 91 30.10 0.902 0.339 6.5 30.15 0.903 0.339 26.8
+LoD 6 103 30.10 0.902 0.338 7.3 30.16 0.903 0.337 30.2
+LoD 7 116 30.11 0.902 0.337 8.2 30.16 0.903 0.337 34.0
+10 Additional Results
+10.1 Comparison with SOTA
+We provide a more detailed comparison with recent state-of-the-art methods.
+We evaluated our method against HAC [7], RDO [37], LightGaussian [11], SOG [27], Compact-3DGS [31], Reduced-3DGS [33], EAGLES [15] and Comp-GS [22], all of which represent the latest advancements in the fields, as also highlighted by  [3].
+All methods were tested on the same hardware setup.
+We observed that all results reported in the respective papers are reproducible.
+However, there are some discrepancies in the LPIPS values, as these are often computed without normalizing pixel values between $[-1,1]$ (which tends to yield higher values, as noted by [5]).
+We also derived rate/distortion curves for some of the methods.
+We use the hyperparameters reported from the respective paper, defining 3 levels of compression, high, medium and low, as shown in Tab.
+7.
+Table 7: Hyperparameters used for deriving the rate/distortion curves for the given methods. Notation is in accordance with the respective papers.
+Method Compression Rate
+High Medium Low
+HAC [7] $\lambda=0.01$ $\lambda=0.001$ $\lambda=0.005$
+RDO  [37] $\lambda_{sh}=0.05$, $\lambda_{gs}=0.005$ $\lambda_{sh}=0.02$, $\lambda_{gs}=0.002$ $\lambda_{sh}=0.005$, $\lambda_{gs}=0.0005$
+Reduced-3DGS [33] $\tau_{\text{std}}=0.06,\tau_{\text{cdist}}=8$ $\tau_{\text{std}}=0.04,\tau_{\text{cdist}}=6$ $\tau_{\text{std}}=0.01,\tau_{\text{cdst}}=1$
+Comp-GS [23] $\lambda=0.01$ $\lambda=0.001$ $\lambda=0.005$
+SoG [27] with SH - no SHs
+For SOG, we identify with low and high the version with SHs and without SHs, respectively.
+In Tab.
+8 all the results for each dataset are reported (this is the complete version of Tab.
+1).
+Notably, our method is the only one trained once, meaning that the curves are generated instantly without the need for re-training.
+Table 8: Main results on the 3 dataset (Mip-Nerf360, Tanks&Temples and DeepBlending). All results were tested on a NVIDIA A40.
+Dataset Mip-NeRF360 Tanks&Temples Deep Blending
+[…]
+Ours 3DGS LoD 0 25.28 0.719 0.407 3.8 100 548 21.82 0.770 0.330 3.6 100 627 29.22 0.888 0.372 3.7 100 814
+LoD 1 26.10 0.751 0.371 5.6 155 457 22.62 0.795 0.302 5.0 144 542 29.57 0.896 0.358 5.3 146 717
+LoD 2 26.63 0.776 0.341 8.6 242 391 23.12 0.813 0.278 7.0 208 470 29.70 0.900 0.348 7.4 212 622
+LoD 3 27.05 0.793 0.314 13.3 381 333 23.46 0.826 0.258 9.7 301 394 29.78 0.902 0.339 10.5 309 540
+LoD 4 27.30 0.804 0.293 20.6 603 282 23.65 0.834 0.244 13.6 437 337 29.78 0.904 0.333 14.8 451 465
+LoD 5 27.40 0.809 0.278 32.1 962 236 23.73 0.838 0.235 19.2 638 278 29.78 0.904 0.328 20.7 658 395
+LoD 6 27.43 0.811 0.271 49.0 1546 197 23.75 0.840 0.230 26.6 935 230 29.77 0.905 0.325 28.9 962 327
+LoD 7 27.42 0.811 0.268 73.7 2497 175 23.75 0.841 0.227 36.8 1374 205 29.77 0.905 0.324 39.8 1406 270
+baseline 27.48 0.816 0.256 787.0 3329 148 23.79 0.849 0.213 434.0 1840 186 29.46 0.904 0.314 665.0 2811 167
+Ours 3DGS-MCMC LoD 0 25.40 0.718 0.406 3.7 100 409 21.92 0.769 0.333 3.5 100 510 29.12 0.887 0.376 3.6 100 633
+LoD 1 26.22 0.750 0.370 5.6 155 348 22.69 0.793 0.303 4.9 144 427 29.44 0.894 0.361 5.0 146 537
+LoD 2 26.75 0.775 0.338 8.4 242 299 23.22 0.812 0.278 6.8 208 359 29.59 0.898 0.350 7.0 212 459
+LoD 3 27.12 0.794 0.310 12.9 381 255 23.62 0.826 0.255 9.5 301 307 29.65 0.900 0.342 9.8 310 422
+LoD 4 27.37 0.806 0.288 19.9 603 214 23.85 0.836 0.238 13.3 438 246 29.70 0.902 0.335 13.6 452 356
+LoD 5 27.50 0.814 0.271 30.8 962 176 23.99 0.841 0.226 18.7 640 206 29.72 0.902 0.329 19.1 660 312
+LoD 6 27.55 0.817 0.261 47.3 1546 143 24.04 0.845 0.217 26.6 938 178 29.73 0.903 0.326 26.7 965 267
+LoD 7 27.55 0.819 0.257 72.2 2497 123 24.05 0.845 0.215 37.4 1379 163 29.71 0.902 0.325 37.5 1411 229
+baseline 27.84 0.831 0.239 794.0 3528 87 24.15 0.862 0.187 435.0 1840 133 29.65 0.904 0.309 667.0 2822 164
+Ours Scaffold-GS LoD 0 24.97 0.710 0.409 4.0 50 145 23.13 0.806 0.295 3.9 50 177 29.76 0.897 0.359 3.9 50 184
+LoD 1 25.98 0.743 0.375 5.4 68 135 23.51 0.821 0.273 4.8 61 162 30.02 0.901 0.350 4.5 58 180
+LoD 2 26.68 0.768 0.345 7.4 94 122 23.72 0.832 0.256 5.9 75 150 30.16 0.904 0.344 5.3 68 173
+LoD 3 27.24 0.789 0.318 10.2 129 108 23.86 0.839 0.242 7.3 93 134 30.26 0.906 0.339 6.2 80 168
+LoD 4 27.61 0.803 0.297 14.1 179 94 23.96 0.844 0.231 8.9 114 128 30.31 0.908 0.335 7.2 93 166
+LoD 5 27.77 0.810 0.284 19.6 248 84 24.01 0.848 0.224 11.0 141 116 30.36 0.909 0.332 8.4 109 163
+LoD 6 27.84 0.812 0.278 27.2 346 75 24.04 0.850 0.219 13.5 174 109 30.38 0.909 0.330 9.8 127 159
+LoD 7 27.87 0.813 0.276 37.9 483 69 24.04 0.851 0.218 16.6 215 100 30.38 0.909 0.329 11.4 149 158
+baseline 27.77 0.813 0.270 171.0 555 79 24.03 0.853 0.210 76.0 247 108 30.28 0.909 0.318 53.0 171 179
+Ours Octree-GS LoD 0 24.61 0.689 0.426 3.7 50 123 23.08 0.798 0.342 3.9 50 94 29.67 0.895 0.358 3.6 50 150
+LoD 1 25.55 0.722 0.391 5.2 69 114 23.68 0.819 0.312 5.3 68 86 29.87 0.898 0.352 4.0 56 146
+LoD 2 26.33 0.752 0.357 7.1 96 104 24.01 0.834 0.286 7.1 93 78 29.98 0.900 0.348 4.5 64 146
+LoD 3 27.01 0.780 0.322 10.0 133 92 24.25 0.847 0.262 9.7 127 72 30.04 0.901 0.344 5.1 72 142
+LoD 4 27.49 0.802 0.291 13.9 186 81 24.40 0.856 0.242 13.2 173 65 30.07 0.902 0.341 5.7 81 139
+LoD 5 27.73 0.813 0.271 19.5 262 71 24.50 0.863 0.227 17.9 235 57 30.10 0.902 0.339 6.5 91 140
+LoD 6 27.85 0.818 0.260 27.5 371 64 24.56 0.867 0.218 24.2 321 51 30.10 0.902 0.338 7.3 103 135
+LoD 7 27.87 0.818 0.257 39.0 527 60 24.56 0.868 0.215 32.8 438 45 30.11 0.902 0.337 8.2 116 135
+baseline 27.96 0.821 0.252 157.2 606 69 24.69 0.873 0.204 131.0 503 49 30.09 0.902 0.338 34.6 133 174
+HAC [7] high 27.17 0.798 0.306 11.5 371 74 24.03 0.843 0.237 7.4 261 82 29.92 0.903 0.250 3.9 180 166
+medium 27.53 0.807 0.290 15.2 425 72 24.05 0.846 0.222 7.9 300 85 29.98 0.904 0.340 4.1 187 173
+low 27.78 0.812 0.277 22.0 501 68 24.37 0.853 0.215 11.1 307 88 30.34 0.909 0.329 6.3 196 166
+RDO [37] high 26.45 0.782 0.322 9.7 862 280 23.19 0.827 0.253 5.5 395 370 29.61 0.903 0.331 7.0 531 335
+medium 26.89 0.796 0.298 15.3 1236 227 23.16 0.833 0.239 8.0 598 307 29.68 0.905 0.323 11.0 891 269
+low 27.05 0.801 0.288 23.3 1860 178 23.32 0.839 0.232 11.9 912 257 29.72 0.906 0.318 18.0 1478 200
+Reduced-3DGS [33] high 27.05 0.807 0.272 22.7 1245 247 23.46 0.840 0.228 10.5 558 384 29.63 0.906 0.318 13.6 804 337
+medium 27.19 0.810 0.267 29.3 1436 234 23.55 0.843 0.223 14.0 656 358 29.70 0.907 0.315 18.3 990 304
+low 27.28 0.813 0.264 45.8 1434 240 23.57 0.844 0.221 20.7 648 360 29.69 0.907 0.314 35.3 988 311
+Comp-GS [23] high 26.40 0.778 0.323 9.0 454 233 23.06 0.816 0.278 6.1 270 334 29.22 0.894 0.369 6.0 311 440
+medium 26.71 0.790 0.305 10.9 470 193 23.32 0.828 0.259 7.3 242 311 29.56 0.899 0.351 7.1 271 315
+low 27.28 0.802 0.283 16.4 485 165 23.62 0.837 0.248 9.8 241 286 29.89 0.904 0.336 10.0 246 254
+SOG [28] high 26.56 0.791 0.241 16.7 2150 119 23.15 0.828 0.198 9.3 1207 216 29.12 0.892 0.270 9.3 800 219
+low 27.08 0.799 0.230 40.3 2176 132 23.56 0.837 0.186 22.8 1242 227 29.26 0.894 0.268 17.7 890 236
+EAGLES [15] - 27.13 0.809 0.278 57.1 1290 156 23.20 0.837 0.241 28.9 651 227 29.75 0.910 0.318 52.4 1192 144
+LightGuassian [11] - 27.24 0.810 0.273 51.0 2197 161 23.55 0.839 0.235 28.5 1211 225 29.41 0.904 0.329 43.2 956 348
+Compact-3DGS [22] - 27.02 0.800 0.287 29.1 1429 108 23.41 0.836 0.238 20.9 842 147 29.76 0.905 0.324 23.8 1053 144
+Table 9: Per-scene results on Tanks&Temples.
+Scene Method/Metric Ours 3DGS Ours 3DGS-MCMC Ours Scaffold-GS Ours Octree-GS
+[…]
+Train baseline 22.19 0.815 0.249 1088 257.3 197 22.30 0.828 0.238 1090 258.0 141 22.23 0.820 0.246 196 60.5 107 22.80 0.848 0.233 526 136.0 44
+LoD 7 22.09 0.800 0.274 812 23.0 196 22.26 0.802 0.278 816 22.3 199 22.31 0.818 0.257 171 13.0 105 22.70 0.842 0.246 458 35.4 41
+LoD 6 22.08 0.799 0.280 602 17.6 234 22.26 0.802 0.278 604 17.2 213 22.30 0.817 0.260 143 10.9 114 22.71 0.840 0.251 334 26.1 47
+LoD 5 22.05 0.795 0.287 446 13.6 281 22.23 0.797 0.287 448 13.0 233 22.27 0.814 0.266 120 9.2 124 22.65 0.835 0.262 243 19.1 56
+LoD 4 21.97 0.789 0.297 331 10.5 340 22.18 0.792 0.296 332 10.0 275 22.21 0.809 0.275 101 7.7 135 22.55 0.825 0.279 177 14.0 63
+LoD 3 21.85 0.780 0.309 245 8.0 398 22.09 0.784 0.308 246 7.7 315 22.12 0.802 0.286 85 6.5 146 22.40 0.813 0.300 129 10.2 71
+LoD 2 21.65 0.769 0.323 182 6.2 465 21.91 0.773 0.323 182 5.9 360 22.01 0.794 0.298 71 5.5 156 22.18 0.797 0.323 94 7.5 78
+LoD 1 21.41 0.756 0.339 135 4.7 517 21.63 0.760 0.340 135 4.5 413 21.86 0.782 0.314 60 4.6 168 21.89 0.780 0.349 69 5.5 86
+LoD 0 20.95 0.738 0.361 100 3.6 610 21.17 0.741 0.362 100 3.4 488 21.60 0.766 0.333 50 3.9 181 21.39 0.756 0.378 50 4.0 94
+Truck baseline 25.39 0.882 0.178 2582 610.8 174 26.00 0.895 0.136 2589 612.0 125 25.82 0.885 0.173 298 91.8 110 26.58 0.898 0.174 481 126.0 54
+LoD 7 25.42 0.882 0.179 1937 50.7 214 25.83 0.889 0.153 1942 52.4 128 25.78 0.884 0.179 259 20.2 95 26.41 0.894 0.184 418 30.1 50
+LoD 6 25.42 0.882 0.180 1268 35.6 227 25.83 0.888 0.156 1271 36.0 143 25.79 0.883 0.179 205 16.1 103 26.42 0.894 0.185 309 22.4 54
+LoD 5 25.40 0.881 0.183 831 24.7 274 25.76 0.886 0.164 832 24.4 179 25.76 0.882 0.182 162 12.8 108 26.36 0.892 0.191 228 16.6 59
+LoD 4 25.33 0.878 0.191 544 16.7 334 25.53 0.879 0.180 545 16.5 218 25.72 0.880 0.188 128 10.1 122 26.25 0.887 0.204 168 12.3 66
+LoD 3 25.07 0.871 0.208 356 11.3 390 25.15 0.868 0.202 357 11.3 299 25.60 0.876 0.198 101 8.0 123 26.09 0.881 0.225 124 9.2 72
+LoD 2 24.59 0.857 0.233 233 7.8 474 24.53 0.850 0.233 233 7.7 359 25.43 0.870 0.213 80 6.4 144 25.84 0.871 0.249 92 6.8 79
+LoD 1 23.83 0.834 0.264 153 5.2 568 23.76 0.827 0.267 153 5.3 441 25.15 0.860 0.232 63 5.0 157 25.46 0.859 0.275 68 5.0 86
+LoD 0 22.68 0.803 0.299 100 3.6 644 22.66 0.798 0.305 100 3.5 531 24.66 0.845 0.256 50 4.0 174 24.77 0.840 0.305 50 3.7 94
+Table 10: Per-scene results on DeepBlending.
+Scene Method/Metric Ours 3DGS Ours 3DGS-MCMC Ours Scaffold-GS Ours Octree-GS
+[…]
+DrJohnson baseline 29.03 0.900 0.318 3298 780.1 146 28.92 0.892 0.320 3308 782.0 173 29.75 0.907 0.322 192 59.2 167 29.20 0.894 0.350 140 162.4 140
+LoD 7 29.28 0.899 0.332 1649 47.9 316 29.00 0.892 0.338 1654 45.1 290 29.76 0.91 0.33 167 12.6 148 29.27 0.894 0.348 122 8.6 130
+LoD 6 29.27 0.899 0.333 1105 34.1 389 29.00 0.892 0.340 1108 31.5 335 29.73 0.91 0.33 141 10.6 151 29.25 0.894 0.349 107 7.6 130
+LoD 5 29.28 0.898 0.337 740 23.9 482 28.98 0.892 0.344 742 21.9 398 29.70 0.905 0.337 118 9.0 154 29.25 0.894 0.350 95 6.7 134
+LoD 4 29.26 0.897 0.342 496 16.6 570 28.97 0.891 0.350 497 15.2 457 29.64 0.903 0.340 100 7.6 158 29.23 0.893 0.352 83 5.9 133
+LoD 3 29.25 0.895 0.350 332 11.4 679 28.87 0.889 0.358 333 10.8 553 29.57 0.901 0.346 84 6.4 163 29.21 0.892 0.355 73 5.2 136
+LoD 2 29.15 0.891 0.361 223 7.9 767 28.77 0.885 0.368 223 7.4 593 29.45 0.898 0.352 71 5.4 171 29.15 0.891 0.359 65 4.6 140
+LoD 1 28.98 0.886 0.374 149 5.4 867 28.58 0.880 0.381 149 5.2 681 29.28 0.894 0.360 59 4.5 176 29.05 0.89 0.36 57 4.1 142
+LoD 0 28.56 0.875 0.391 100 3.7 970 28.17 0.871 0.399 100 3.6 818 28.97 0.888 0.370 50 3.8 179 28.85 0.89 0.37 50 3.6 144
+Playroom baseline 29.89 0.907 0.309 2324 550 187 30.39 0.916 0.298 2335 552.0 154 30.81 0.912 0.314 150 46.2 190 30.99 0.911 0.326 126 32.6 186
+LoD 7 30.27 0.911 0.316 1162 31.7 224 30.43 0.913 0.311 1168 30.0 168 30.99 0.912 0.325 130 10.2 168 30.95 0.911 0.327 110 7.8 141
+LoD 6 30.27 0.911 0.317 818 23.7 266 30.46 0.913 0.311 822 21.9 198 31.02 0.912 0.326 114 8.9 166 30.95 0.911 0.327 98 6.9 141
+LoD 5 30.28 0.911 0.320 577 17.5 308 30.45 0.913 0.315 579 16.2 225 31.01 0.912 0.327 99 7.8 172 30.94 0.911 0.329 88 6.2 145
+LoD 4 30.29 0.911 0.324 406 13.0 359 30.43 0.913 0.320 407 12.0 255 30.98 0.912 0.329 86 6.8 174 30.91 0.910 0.330 78 5.6 145
+LoD 3 30.30 0.910 0.328 286 9.5 402 30.43 0.912 0.325 287 8.8 290 30.94 0.911 0.332 75 6.0 173 30.87 0.910 0.333 70 5.0 149
+LoD 2 30.25 0.909 0.334 202 7.0 477 30.40 0.911 0.333 202 6.6 326 30.87 0.910 0.336 66 5.2 175 30.82 0.909 0.336 63 4.5 151
+LoD 1 30.15 0.906 0.342 142 5.1 566 30.31 0.908 0.341 142 4.8 392 30.77 0.908 0.340 57 4.5 183 30.70 0.907 0.340 56 4.0 149
+LoD 0 29.89 0.901 0.354 100 3.8 658 30.07 0.904 0.353 100 3.6 449 30.56 0.905 0.347 50 4.0 189 30.50 0.905 0.346 50 3.6 156
+10.2 LoD Visualizations
+In Fig.
+101112 we show all LoD levels for each dataset on 3DGS.
+We observe that the transition between levels is smooth, with each level progressively adding more detail without introducing artifacts.
+We observe that lower levels exhibit a loss of detail in high-frequency regions, particularly in background areas.
+Examples include the trees in the bicycle scene, the details of the building in the background of the garden scene, and the reflections and transparencies in the background of the counter scene.
+Figure 10: Full LoD visualization with 3DGS on Tanks&Temples dataset.
+Figure 11: Full LoD visualization with 3DGS on DeepBlending dataset.
+Figure 12: Full LoD visualization with 3DGS on Mip-NeRF360 dataset.
+10.3 Per Scene Results
+In Tab. 11910 we provide per-scene results for all 3 tested models across every LoD level with compression, along with the corresponding baseline values.
+Table 11: Per-scene results on Mip-NeRF360.
+Scene Method/Metric Ours 3DGS Ours 3DGS-MCMC Ours Scaffold-GS Ours Octree-GS
+[…]
+Bicycle baseline 25.17 0.765 0.24 6091 1371 89 25.53 0.791 0.211 6091 1371 60 25.18 0.747 0.289 879 749.0 60 25.31 0.761 0.256 1049 273.8 52
+LoD 7 25.04 0.757 0.254 4569 134.0 141 25.24 0.770 0.236 4569 130.0 110 25.27 0.748 0.291 764 60.3 51 25.16 0.758 0.260 912 67.8 45
+LoD 6 25.04 0.756 0.258 2646 84.3 170 25.23 0.770 0.240 2646 81.2 131 25.25 0.748 0.293 518 40.8 54 25.18 0.757 0.263 602 44.9 47
+LoD 5 25.00 0.752 0.273 1533 51.3 226 25.16 0.760 0.261 1533 48.9 177 25.22 0.745 0.300 351 27.9 61 25.09 0.749 0.279 398 29.9 53
+LoD 4 24.89 0.742 0.297 888 30.3 286 24.98 0.750 0.290 888 29.1 232 25.01 0.728 0.326 238 18.9 75 24.74 0.721 0.320 263 19.9 63
+LoD 3 24.68 0.724 0.327 514 17.8 353 24.68 0.720 0.326 514 17.3 287 24.67 0.703 0.357 161 12.9 87 24.20 0.675 0.374 174 13.2 74
+LoD 2 24.34 0.694 0.365 298 10.4 425 24.32 0.690 0.370 298 10.3 357 24.20 0.669 0.393 109 8.8 101 23.69 0.630 0.421 115 8.7 89
+LoD 1 23.85 0.653 0.411 173 6.1 503 23.81 0.650 0.417 173 6.1 421 23.59 0.629 0.430 74 5.9 113 23.17 0.591 0.460 76 5.8 102
+LoD 0 23.11 0.599 0.461 100 3.6 625 23.11 0.590 0.466 100 3.6 478 22.69 0.582 0.470 50 4.0 125 22.50 0.551 0.496 50 3.8 110
+Bonsai baseline 32.20 0.942 0.253 1258 283 215 32.32 0.943 0.261 1258 283 106 32.80 0.949 0.233 400 524.0 75 32.79 0.949 0.236 240 62.4 106
+LoD 7 31.89 0.939 0.266 944 25.8 211 31.92 0.940 0.265 944 25.6 154 32.95 0.950 0.242 348 25.5 69 32.72 0.948 0.238 209 14.2 89
+LoD 6 31.89 0.939 0.266 685 19.6 225 31.91 0.940 0.265 685 18.8 159 32.90 0.950 0.243 264 19.5 74 32.71 0.947 0.240 170 11.7 92
+LoD 5 31.83 0.938 0.268 497 14.8 245 31.87 0.940 0.270 497 14.3 182 32.82 0.948 0.246 200 14.8 86 32.60 0.946 0.243 139 9.6 99
+LoD 4 31.77 0.937 0.273 361 11.3 276 31.76 0.940 0.277 361 10.8 207 32.58 0.945 0.252 152 11.3 100 32.36 0.942 0.250 113 7.8 105
+LoD 3 31.67 0.935 0.281 262 8.6 296 31.59 0.930 0.286 262 8.2 230 32.09 0.939 0.263 115 8.6 113 31.80 0.935 0.262 92 6.4 112
+LoD 2 31.39 0.930 0.292 190 6.4 338 31.28 0.930 0.299 190 6.2 259 31.26 0.929 0.281 87 6.5 128 30.83 0.924 0.280 75 5.2 119
+LoD 1 30.85 0.922 0.307 138 4.8 379 30.73 0.920 0.316 138 4.8 290 30.18 0.913 0.304 66 4.9 142 29.58 0.908 0.304 61 4.3 124
+LoD 0 29.69 0.906 0.330 100 3.7 434 29.54 0.900 0.339 100 3.6 311 28.44 0.887 0.338 50 3.8 154 28.18 0.887 0.331 50 3.5 135
+Counter baseline 29.01 0.909 0.258 1205 271 150 28.95 0.911 0.257 1205 271 82 29.67 0.918 0.242 273 383.0 73 30.20 0.925 0.232 336 86.4 63
+LoD 7 28.86 0.904 0.271 904 25.2 158 28.80 0.900 0.269 904 24.9 119 29.85 0.919 0.249 238 17.3 70 30.08 0.924 0.236 292 19.5 56
+LoD 6 28.87 0.904 0.272 660 19.3 170 28.80 0.900 0.270 660 18.8 128 29.80 0.918 0.252 190 14.0 75 30.04 0.922 0.242 227 15.3 60
+LoD 5 28.86 0.904 0.274 482 15.0 186 28.79 0.900 0.273 482 14.5 148 29.72 0.915 0.260 152 11.3 82 29.91 0.917 0.254 176 12.0 68
+LoD 4 28.83 0.902 0.279 352 11.5 206 28.73 0.900 0.279 352 11.3 163 29.55 0.909 0.272 122 9.0 89 29.68 0.908 0.271 137 9.3 74
+LoD 3 28.74 0.899 0.288 257 8.8 221 28.62 0.900 0.288 257 8.6 179 29.26 0.900 0.288 97 7.2 96 29.28 0.896 0.292 107 7.3 81
+LoD 2 28.54 0.893 0.299 188 6.6 247 28.45 0.890 0.299 188 6.6 196 28.79 0.887 0.308 78 5.8 103 28.67 0.881 0.315 83 5.7 87
+LoD 1 28.21 0.885 0.314 137 5.0 281 28.19 0.880 0.314 137 5.0 225 28.21 0.872 0.330 62 4.7 111 27.97 0.864 0.339 64 4.4 94
+LoD 0 27.61 0.871 0.334 100 3.8 329 27.61 0.870 0.334 100 3.8 255 27.43 0.853 0.354 50 3.7 119 27.09 0.844 0.365 50 3.4 103
+Flowers baseline 21.54 0.606 0.366 3598 810 178 21.96 0.638 0.324 3598 810 77 21.25 0.577 0.407 669 592.0 90 21.52 0.599 0.375 777 202.4 70
+LoD 7 21.44 0.596 0.378 2698 80.7 231 21.65 0.610 0.350 2698 82.2 118 21.27 0.579 0.408 582 47.6 80 21.51 0.598 0.377 676 52.8 61
+LoD 6 21.44 0.595 0.382 1685 54.7 252 21.62 0.610 0.362 1685 53.1 151 21.27 0.579 0.410 410 33.7 88 21.51 0.597 0.379 466 36.6 67
+LoD 5 21.42 0.592 0.392 1053 36.0 305 21.54 0.600 0.376 1053 35.4 199 21.27 0.577 0.414 289 23.9 97 21.49 0.594 0.386 321 25.5 74
+LoD 4 21.35 0.584 0.408 657 23.1 358 21.36 0.580 0.398 657 22.5 248 21.21 0.572 0.422 203 16.9 107 21.37 0.585 0.399 221 17.7 84
+LoD 3 21.19 0.569 0.428 411 14.8 438 21.10 0.560 0.423 411 14.4 304 20.93 0.556 0.439 143 12.0 125 21.02 0.563 0.421 153 12.3 99
+LoD 2 20.89 0.543 0.453 256 9.5 496 20.77 0.530 0.450 256 9.3 358 20.51 0.532 0.459 101 8.5 143 20.33 0.528 0.451 105 8.5 118
+LoD 1 20.44 0.510 0.480 160 6.0 565 20.29 0.490 0.479 160 5.9 430 20.02 0.504 0.481 71 6.0 160 19.55 0.487 0.483 73 5.9 137
+LoD 0 19.76 0.470 0.511 100 3.9 703 19.58 0.450 0.513 100 3.8 493 19.38 0.470 0.506 50 4.2 173 18.75 0.444 0.517 50 4.1 146
+Garden baseline 27.39 0.867 0.123 5872 1322 100 27.68 0.874 0.114 5872 1322 76 27.28 0.850 0.157 732 $1$.0 58 27.61 0.860 0.134 856 220.0 44
+LoD 7 27.21 0.859 0.139 4404 133.2 126 27.31 0.860 0.135 4404 126.4 107 27.19 0.845 0.167 636 50.0 52 27.52 0.854 0.147 744 57.0 39
+LoD 6 27.20 0.859 0.142 2564 85.2 160 27.29 0.860 0.139 2564 80.4 141 27.15 0.844 0.172 443 35.1 60 27.42 0.850 0.154 506 39.1 45
+LoD 5 27.11 0.854 0.156 1493 52.7 218 27.16 0.850 0.155 1493 50.0 193 26.94 0.835 0.188 308 24.6 73 27.12 0.837 0.178 344 26.7 56
+LoD 4 26.79 0.840 0.187 870 31.5 284 26.78 0.840 0.188 870 30.3 249 26.58 0.817 0.219 214 17.2 91 26.66 0.813 0.219 234 18.2 71
+LoD 3 26.02 0.807 0.241 506 18.6 355 26.11 0.810 0.240 506 18.2 309 25.97 0.788 0.262 149 12.0 109 25.94 0.775 0.277 159 12.4 85
+LoD 2 24.95 0.755 0.308 295 11.0 447 25.09 0.760 0.305 295 10.6 374 25.06 0.745 0.315 103 8.4 130 24.92 0.723 0.336 108 8.5 103
+LoD 1 23.84 0.691 0.373 172 6.4 543 23.98 0.700 0.371 172 6.3 449 24.06 0.694 0.368 72 5.8 149 23.85 0.664 0.391 74 5.8 115
+LoD 0 22.66 0.619 0.436 100 3.9 682 22.83 0.620 0.436 100 3.7 552 22.71 0.633 0.421 50 4.1 164 22.55 0.598 0.442 50 3.9 124
+Kitchen baseline 31.51 0.928 0.155 1827 411 122 31.83 0.929 0.162 1827 411 76 31.81 0.932 0.146 332 879.0 57 31.99 0.934 0.145 319 82.3 55
+LoD 7 31.40 0.924 0.166 1371 38.6 136 31.16 0.920 0.176 1371 36.1 111 31.81 0.931 0.153 289 22.2 54 31.58 0.931 0.152 278 19.7 51
+LoD 6 31.43 0.924 0.167 943 28.1 161 31.15 0.920 0.177 943 25.9 128 31.72 0.930 0.156 225 17.4 59 31.49 0.929 0.156 217 15.5 56
+LoD 5 31.38 0.922 0.172 649 20.6 195 31.08 0.920 0.181 649 19.0 151 31.54 0.927 0.162 175 13.5 69 31.27 0.925 0.165 170 12.1 64
+LoD 4 31.09 0.918 0.183 446 14.9 230 30.92 0.920 0.188 446 13.7 167 31.23 0.922 0.174 136 10.6 79 30.94 0.919 0.178 133 9.5 72
+LoD 3 30.40 0.909 0.201 307 10.8 271 30.70 0.910 0.197 307 10.1 197 30.72 0.913 0.191 106 8.3 90 30.29 0.908 0.198 104 7.5 80
+LoD 2 29.45 0.898 0.221 211 7.8 320 30.24 0.910 0.211 211 7.4 211 29.98 0.900 0.215 83 6.4 100 29.39 0.892 0.226 82 5.9 87
+LoD 1 28.58 0.884 0.245 145 5.6 356 29.59 0.900 0.230 145 5.3 227 29.06 0.883 0.243 64 5.0 111 28.46 0.874 0.255 64 4.6 95
+LoD 0 27.36 0.864 0.275 100 4.0 411 28.38 0.880 0.258 100 3.9 277 27.62 0.859 0.278 50 3.9 119 27.19 0.849 0.291 50 3.6 101
+Room baseline 31.48 0.919 0.287 1539 1088 154 31.96 0.928 0.268 4833 1088 130 32.20 0.931 0.260 272 267.0 105 32.65 0.937 0.248 334 86.4 83
+LoD 7 31.65 0.919 0.298 1154 29.0 155 31.84 0.920 0.288 1154 29.4 143 32.46 0.932 0.269 236 17.2 91 32.59 0.935 0.254 291 21.7 69
+LoD 6 31.65 0.919 0.298 814 21.2 167 31.84 0.920 0.288 814 21.1 148 32.45 0.932 0.270 189 13.9 92 32.57 0.934 0.256 226 17.0 73
+LoD 5 31.64 0.918 0.300 574 15.6 190 31.83 0.920 0.289 574 14.7 154 32.41 0.932 0.272 152 11.2 104 32.46 0.932 0.262 176 13.2 77
+LoD 4 31.60 0.917 0.305 405 11.5 214 31.78 0.920 0.295 405 10.8 172 32.31 0.929 0.279 122 9.0 106 32.28 0.928 0.273 137 10.3 83
+LoD 3 31.49 0.915 0.314 285 8.6 239 31.61 0.920 0.304 285 7.9 191 32.08 0.925 0.290 97 7.2 116 32.00 0.922 0.290 106 8.0 89
+LoD 2 31.33 0.911 0.324 201 6.3 271 31.49 0.910 0.315 201 5.8 207 31.68 0.918 0.307 78 5.8 124 31.50 0.913 0.310 83 6.3 95
+LoD 1 31.08 0.906 0.337 142 4.6 299 31.20 0.910 0.329 142 4.4 241 31.15 0.909 0.325 62 4.6 126 30.88 0.903 0.331 64 4.9 99
+LoD 0 30.62 0.898 0.354 100 3.3 351 30.73 0.900 0.348 100 3.2 276 30.39 0.898 0.346 50 3.7 131 30.05 0.891 0.353 50 3.8 105
+Stump baseline 26.56 0.771 0.250 4833 842 166 27.35 0.806 0.217 3740 842 74 26.65 0.764 0.300 734 642.0 104 26.75 0.770 0.283 642 167.8 87
+LoD 7 26.78 0.775 0.256 3624 109.8 214 27.19 0.800 0.234 3624 107.1 105 26.82 0.766 0.300 638 52.8 82 26.77 0.769 0.289 558 42.0 73
+LoD 6 26.79 0.775 0.257 2170 70.9 232 27.18 0.790 0.240 2170 67.9 132 26.81 0.766 0.300 443 36.6 87 26.78 0.769 0.290 395 29.8 77
+LoD 5 26.79 0.775 0.264 1299 44.6 279 27.18 0.790 0.252 1299 43.1 181 26.81 0.766 0.301 308 25.4 92 26.73 0.766 0.297 280 21.2 84
+LoD 4 26.76 0.772 0.279 778 27.5 352 27.08 0.790 0.273 778 26.8 238 26.76 0.763 0.309 214 17.7 103 26.47 0.753 0.319 198 15.1 97
+LoD 3 26.64 0.765 0.303 466 16.8 434 26.88 0.770 0.301 466 16.6 302 26.37 0.746 0.332 149 12.3 124 25.96 0.727 0.353 141 10.8 113
+LoD 2 26.33 0.747 0.338 279 10.3 528 26.50 0.750 0.338 279 10.3 368 25.75 0.720 0.365 103 8.6 147 25.30 0.693 0.394 100 7.7 129
+LoD 1 25.80 0.717 0.380 167 6.4 668 25.89 0.720 0.382 167 6.3 428 24.97 0.687 0.402 72 6.0 166 24.52 0.655 0.435 71 5.5 144
+LoD 0 24.89 0.672 0.431 100 3.9 778 24.96 0.670 0.434 100 3.9 527 23.94 0.645 0.444 50 4.2 177 23.62 0.613 0.476 50 3.9 151
+Treehill baseline 22.47 0.634 0.375 3740 749 160 22.94 0.655 0.339 3329 749 98 23.08 0.644 0.401 704 662.0 87 22.85 0.651 0.358 897 233.0 61
+LoD 7 22.52 0.626 0.388 2805 87.3 204 22.85 0.640 0.358 2805 87.7 138 23.20 0.646 0.403 612 48.0 74 22.90 0.652 0.361 780 56.1 54
+LoD 6 22.56 0.626 0.395 1742 58.0 234 22.94 0.640 0.368 1742 58.2 165 23.22 0.646 0.404 428 33.5 82 22.93 0.652 0.365 527 37.9 59
+LoD 5 22.58 0.626 0.406 1082 37.6 284 22.93 0.640 0.382 1082 37.4 203 23.22 0.645 0.408 299 23.5 89 22.95 0.651 0.373 356 25.7 66
+LoD 4 22.61 0.624 0.423 672 24.1 331 22.90 0.630 0.403 672 23.9 251 23.23 0.644 0.416 209 16.4 96 22.88 0.644 0.393 240 17.4 78
+LoD 3 22.59 0.619 0.443 417 14.8 387 22.79 0.620 0.428 417 14.7 299 23.10 0.634 0.436 146 11.5 111 22.65 0.622 0.430 162 11.8 93
+LoD 2 22.49 0.609 0.466 259 9.4 450 22.61 0.610 0.457 259 9.3 355 22.85 0.614 0.464 102 8.0 120 22.32 0.589 0.477 110 8.0 106
+LoD 1 22.25 0.594 0.496 161 5.8 519 22.28 0.590 0.491 161 6.0 418 22.56 0.592 0.494 72 5.6 135 21.97 0.556 0.520 74 5.4 122
+LoD 0 21.85 0.571 0.531 100 3.7 616 21.83 0.570 0.529 100 3.8 508 22.12 0.564 0.527 50 3.9 146 21.56 0.525 0.559 50 3.6 133